### v1

# RobustBlack: Challenging Black-Box Adversarial Attacks on State-of-the-Art Defenses

Abstract
Although adversarial robustness has been extensively studied in white-box settings, recent advances in black-box attacks (including transfer- and query-based approaches) are primarily benchmarked against weak defenses, leaving a significant gap in the evaluation of their effectiveness against more recent and moderate robust models (e.g., those featured in the Robustbench leaderboard). In this paper, we question this lack of attention from black-box attacks to robust models. We establish a framework to evaluate the effectiveness of recent black-box attacks against both top-performing and standard defense mechanisms, on the ImageNet dataset. Our empirical evaluation reveals the following key findings: (1) the most advanced black-box attacks struggle to succeed even against simple adversarially trained models; (2) robust models that are optimized to withstand strong white-box attacks, such as AutoAttack, also exhibits enhanced resilience against black-box attacks; and (3) robustness alignment between the surrogate models and the target model plays a key factor in the success rate of transfer-based attacks.
1 Introduction
Since the discovery that deep learning models are susceptible to minor input disturbances, yielding adversarial examples (AE) (Goodfellow et al., 2015), the development of robust models has become one of the most active topics in the machine learning community. This topic has been thoroughly explored for white-box settings, leading to standardized attacks (Croce & Hein, 2020b) and benchmarks (Croce et al., 2020).
However, realistic evaluations of adversarial attacks remain less explored. It involves scenarios where the attacker has limited or no knowledge of the target model to be attacked. These scenarios are referred to as black-box (or gray-box with partial knowledge (Guo et al., 2018)). Black-box attacks belong to one of the two following settings: When the output of the model provides both category labels and scores, it is referred to as a score-based black-box attack. Conversely, if the output includes only category labels, it is named a decision-based black-box attack (Brendel et al., 2018). Given that most computer vision Application Programming Interface (APIs) (Google, 2023; Imagga, 2023) generally provide image prediction categories alongside scores, most research in a black-box setting considers score-based black-box attacks. Hence, this work focuses on this category.
State of the art (SOTA) black-box attacks leverage one or both of the following mechanisms: (1) adversarial examples transferability and (2) iterative queries with meta-heuristics. Transferability methods craft adversarial examples on a specifically built surrogate model (with white-box access) such that this example should transfer to (i.e. also fools) a target model (to which they have black-box access). Iterative query methods use query feedback from the target model to guide or refine the optimization process to uncover its vulnerabilities.
Transferability is difficult to achieve when the target differs from the surrogate (in terms of architecture, training, or dataset) and the factors behind transferability (or lack thereof) remain an active field of study (Charles et al., 2020; Dong et al., 2018a; Li et al., 2018; Wu et al., 2020a). This is because adversarial attacks crafted on the surrogate lead to examples that maximize the loss function of the surrogate model (Goodfellow et al., 2015; Kurakin et al., 2017), while the target model has a different loss function. Methods to improve transferability mostly rely on building diversity during the optimization of adversarial examples (Li et al., 2018; Wu et al., 2020a; Xie et al., 2019a).
Iterative-query attacks also face multiple challenges: With increased search space (large images, for example), common query attacks show a significant decrease in the Attack Success Rate (ASR) or require large amounts of queries to achieve sufficient ASR (Mohaghegh Dolatabadi et al., 2020; Huang & Zhang, 2020; Feng et al., 2022). BASES (Cai et al., 2022b), a recent Black-box Surrogate Ensemble Search attack achieved more than 95% ASR with a few queries (<5) on average across a large scope of non-robust models.
Though the initial motivation for black-box attacks and their evaluation protocol is to improve the realism of robustness evaluations, they generally suffer from a common pitfall: they do not consider robust models which have already demonstrated their improved performance against white-box attacks. We hypothesize that the evaluation results of black-box attacks can be misleading because even simple robustification mechanisms can be sufficient to successfully evade black-box attacks. Going further, we want to assess whether innovations in defenses provide positive benefits against black-box attacks like they did in white-box settings.
To the best of our knowledge, this paper is the first to study the effectiveness of black-box attacks against strandardized defenses to demonstrate the need to confront black-box attacks to existing defenses.
Our contributions can be summarized as follows:
  1. 1.
    We demonstrate that simple adversarial training mechanisms reduce the effectiveness of black-box attacks proven effective against standard models, underscoring the need for more advanced black-box attack strategies to address the robustness of real-world models and systems.
  2. 2.
    We show that white-box robustness could be served as a proxy for black-box robustness. Defenses optimized against AutoAttack generalize well to black-box scenarios.
  3. 3.
    We demonstrate that these effective defense mechanisms can inadvertently contribute to enhancing the success rate of black-box attacks. By using robust models as surrogates, attacks can generate adversarial examples more likely to transfer to robust models, leading to an average increase in success rate of 6.42 percentage points across attacks and target models.
Our contributions send a clear message to the adversarial research community regarding the importance of considering white-box defenses in the evaluation of black-box attacks. They also urge the researchers working on defense mechanisms to study how their defenses can help improve transfer-based attacks.
2 Related Work
2.1 Black-Box Attacks: Transfer-Based and Query-Based Methods
Black-box attacks have seen notable advancements, especially within transfer-based and query-based approaches. Transfer-based attacks rely on adversarial examples generated by a surrogate model, which are then tested on a target model. Early approaches like Projected gradient descent (PGD) (Madry et al., 2017) laid the groundwork for such attacks by iteratively optimizing perturbations to maximize the success rate on the target model . This foundational method was enhanced by the Momentum Iterative Fast Gradient Sign Method (MI-FGSM), which incorporated momentum into PGD to improve transferability and stability across iterations (Dong et al., 2018b). Further innovations, such as the Diverse Input Fast Gradient Sign Method (DI-FGSM) and Translation-Invariant FGSM (TI-FGSM), introduced input transformations and image translations, respectively, to increase the robustness and generalizability of adversarial examples across different models (Xie et al., 2019b; Dong et al., 2019).
Subsequent methods like Variance Tuning (VMI, VNI) refined gradient calculations by considering gradient variance across iterations (Wang & He, 2021), leading to more effective adversarial perturbations. Similarly, ADMIX leveraged diverse inputs by mixing the target image with randomly sampled images to generate more transferable attacks (Wang et al., 2021). Skip Gradient Method (SGM) focused on model architecture, by using skip connections which facilitate the creation of adversarial examples that have high transferability (Wu et al., 2020b). Universal Adversarial Perturbations (UAP) took a different approach by focusing on universality, generating a single small perturbation that can effectively disrupt a classifier across most natural images (Moosavi-Dezfooli et al., 2017). Notably, Ghost networks (GHOST) and Large Geometric Vicinity (LGV) contributed to transferability by modifying skip connections in surrogate models and by exploring variations in surrogate models vicinity, respectively (Li et al., 2020; Gubri et al., 2022).
In query-based attacks, which generally have higher success rates than transfer attacks, methods like Zeroth Order Optimization (ZOO), Decision Boundary Attack (DBA), Natural Evolutionary Strategies (NES),and HopSkipJump rely on iterative query feedback to optimize perturbations, often not utilizing surrogate models to boost efficiency (Brendel et al., 2018; Chen et al., 2017; Ilyas et al., 2018; Chen et al., 2020). Recent work, such as BASES and Transferable Model-based Embedding (TREMBA) (Cai et al., 2022a; Huang & Zhang, 2019), integrates both surrogate representations and query feedback to create highly effective black-box attacks. BASES leverages an ensemble of surrogates with query feedback to dynamically adjust surrogate weighting, while TREMBA utilizes a generator trained with surrogate models, exploring latent space for more effective query attacks.
2.2 Defenses Against AutoAttack
In parallel, adversarial defenses have evolved to mitigate the risks posed by increasingly sophisticated white-box attacks. As defenses became more robust, the need for a reliable, standardized benchmark to accurately assess their effectiveness grew increasingly apparent. In response, AutoAttack emerged as a reliable, parameter-free method for evaluating adversarial robustness, offering a computationally affordable, and a standardized benchmark applicable to various models (Croce & Hein, 2020b). Building on this, Robustbench established a leaderboard specifically for evaluating adversarial robustness, as outlined by (Croce et al., 2020). This leaderboard includes results on the ImageNet large-scale datasets (Deng et al., 2009), and offers a ranking comparison of various defense strategies across specific perturbation budgets. Among the defenses featured on the leaderboard are those proposed by (Salman et al., 2020; Singh et al., 2024; Liu et al., 2024; Bai et al., 2024; Madry et al., 2017), offering a thorough assesment on the effectiveness of the leading defense mechanisms. Traditional defenses, such as Madry’s adversarial training, initially combined clean and adversarial data during training, establishing a baseline for robustness (Madry et al., 2017). Recent advancements have incorporated complex pre-training schemes and data augmentations (e.g., ConvStem architectures, RandAugment, MixUp, and CutMix). Singh et al. (2024) made alterations to the convnets architecture, particularly substituting PatchStem with ConvStem, and modified the training scheme to enhance the robustness against previously unencountered l1 and l2 threat models. Liu et al. (2024) further advances the training scheme with large-scale pre-training, combined with label smoothing, weight decay, and Exponential Moving Averaging (EMA) to enhance generalization against unseen adversarial examples. Bai et al. (2024) builds on a giving defense framework (Liu et al., 2024), focusing specifically on balancing clean and robust accuracy through the nonlinear mixing of robust and standard models, addressing the often-seen tradeoff between clean and robust performance.
2.3 Robustness Under Black-Box Setting
In this part, we focus on how our work introduces a novel contribution compared to previous studies on defenses against black-box adversarial attacks in image classification tasks. We discuss four representative works (Dong et al., 2020; Mahmood et al., 2021a; b; Ghaffari Laleh et al., 2022), highlighting their key contributions and how our approach builds upon or complements these efforts by addressing gaps and introducing new perspectives in this domain.
Dong et al. (2020) evaluated adversarial robustness in image classification tasks against white-box and black-box attacks. Their work addressed various defense techniques such as robust training, input transformation, randomization, certified defenses, and model ensembling. Our study seeks to complement this by focusing specifically on SOTA robust training-based defenses. Moreover, our work explores a broader landscape of black-box attacks by including recent methods such as ADMIX, BASES, and TREMBA. In doing so, we provide an analysis of leading robust training defenses against recent black-box adversarial strategies.
Mahmood et al. (2021a) noted that the majority of existing defenses primarily address white-box attacks, neglecting the crucial aspect of black-box adversarial robustness. They provided a wide evaluation of adversarial defenses with a Convolutional neural network (CNN) architecture on benchmark datasets such as CIFAR-10 (Krizhevsky, 2009) and Fashion-MNIST (Xiao et al., 2017) against 12 attacks. In a follow-up work, Mahmood et al. (2021b) extended this investigation to Vision Transformers (ViTs), evaluating their adversarial robustness against white-box attacks and two black-box attacks on ImageNet. They suggested that ensemble defenses can enhance robustness when attackers lack access to model gradients. Additionally, They examined the transferability of adversarial examples between CNNs and transformers. Building upon and complementing these efforts, we conduct our evaluations on ImageNet, incorporating both CNN and transformer architectures, and including an ensemble defense strategy that combines robust and standard base models. Our approach further investigates transferability between adversarially trained and standardly trained models. We leverage SOTA robust models to challenge the attacks used to evaluate robustness.
In the field of computational pathology, Ghaffari Laleh et al. (2022) examined adversarial robustness with a focus on CNN and transformer models. This work addressed certain adversarial attacks (including Fast Adaptive Boundary (FAB) (Croce & Hein, 2020a), and Square (Andriushchenko et al., 2020)), as well as adversarial training using (PGD) with Dual-Batch Normalization (DBN). Our work extends this line of inquiry by incorporating SOTA adversarially trained models and a more comprehensive suite of black-box attacks, providing additional context for these attacks by increasing their number and offering a more challenging confrontation against SOTA defenses.
3 Success Rate of Black-Box Attacks Against Adversarial Training
As preliminaries, we investigate the effectiveness of 12 black-box adversarial attacks against Madry adversarial training (Madry et al., 2017) on a single, relatively simple ResNet50 model (He et al., 2016). The purpose of these experiments is to check to what extent black-box attacks, typically successful against vanilla models, remain as successful when confronted to adversarially trained models. This preliminary experiment provides us with the first insight into the impact of defenses studied in white-box settings on the effectiveness of black-box attacks.
3.1 Experimental Setup
We evaluated the attack success rate of 11 black-box adversarial attacks against a standardly trained model (vanilla) and an adversarially trained model. We selected a ResNet50 vanilla target and a ResNet50 robust target that underwent adversarial training. We opted for the same model architecture to eliminate the variations that could arise from differing model architectures. We followed the Robustbench evaluation protocol (Croce et al., 2020) and reused the original implementation of each attack. All attacks were untargeted and bound to an $L_{\infty}$ maximum perturbation with a distance of $4/255$. We evaluated the ASR on 5000 examples from the ImageNet validation set, we used the same image identifiers (IDs) as in RobustBench. All the experiments are repeated with three randoms seeds.
As outlined in 2.1, we selected nine transfer-based and two query-based adversarial attacks to provide a comprehensive evaluation of adversarial techniques, balancing foundational, and SOTA approaches. For transfer attacks, our selection emphasizes the diversity of techniques by including methods targeting input transformations (TI, DI), gradient optimization (MI, VNI, VMI, ADMIX), universality (UAP), architectural nuances (GHOST), and geometry (LGV). For query-based attacks, we focused on methods that improve query efficiency and performance by leveraging surrogate models, and that rely on different optimization techniques. BASES uses a small dimensional search space by modifying the weights given to each surrogate based on the feedback received from querying the target, while TREMBA learns a low-dimensional embedding then performs an efficient search within this space. As discussed in 2.1, all query attacks in this study require surrogate models. The hyperparameters for all transfer and query attacks are detailed in the Appendix A and B, respectively.
We employed surrogate models that are either individually larger than the target model or collectively larger as an ensemble, allowing us to leverage the full capacity of the attack. For all single-surrogate attacks, including the five FGSM-based baseline attacks, ADMIX, UAP, GHOST, and LGV, the surrogate is a WideResNet50-2 (Zagoruyko, 2016) with standard training. For TREMBA, the surrogates are an ensemble of four models with standard training: VGG16 (Simonyan & Zisserman, 2014), ResNet18, SqueezeNet (Iandola et al., 2016), and GoogLeNet (Szegedy et al., 2015) as provided in their repository. For BASES, the surrogates are an ensemble of 10 models with standard training: VGG16BN, ResNet18, SqueezeNet11, GoogLeNet, MnasNet10 (Tan et al., 2019), DenseNet161 (Huang et al., 2017), EfficientNetB0 (Tan & Le, 2019), RegNetY400MF (Radosavovic et al., 2020), ResNeXt10132x8d (Xie et al., 2017), ConvNeXtSmall (Liu et al., 2022) from . The models used in this study were obtained from the torchvision library (maintainers & contributors, 2016) and the RobustBench benchmark Croce et al. (2020). We excluded the SGM attack from this section because, at the time of this study, its implementation supports only ResNet and DenseNet surrogates, while our setup uses WideResNet50-2.
3.2 Results
We evaluate the ASR of the black-box attacks against standard and adversarially trained Resnet-50 models. We illustrate the mean ASR using bars, and the standard deviation (STD) as error bars in Figure 1. Against the undefended model, all attacks except UAP and LGV achieve more than 40% ASR, with the most recent attacks BASES and TREMBA reaching 81.08% ± 0.49 and 89.56% ± 0.09, respectively. However, all black-box attacks fail against the robustified model with an ASR of less than 4%. The most recent attacks only bring about marginal improvement over the simplest FGSM attacks. The most effective attack in our evaluation, TREMBA achieves 3.56% ASR while TI-FGSM reaches 2.81% ASR. To ensure that the reduced effectiveness of black-box attacks against the adversarially trained ResNet50 model is not due to budget limitation, we quadruple the budget of the best black-box attacks (BASES and TREMBA). Even with this increased budget, both attacks still significantly underperform against the adversarially trained ResNet50 compared to their performance on the vanilla ResNet50 model, as detailed in Appendix C.
Figure 1: The blue bars (with error bars) show the success rates and standard deviations for the vanilla ResNet50 model, while the orange bars (with error bars) show the results for the robust ResNet50 model.
Insight 1 The effectiveness of black-box attacks drastically decreases against a simple adversarially trained model, with the strongest attack seeing its ASR drops from 89.56% to 3.56%.
4 Effectiveness of White-Box Defenses Against Black-Box Attacks
Given that even the most recent advances in black-box attacks remain ineffective against Madry adversarial training, we investigate whether recent innovations in defense mechanisms (studied in white-box settings) further reduce the effectiveness of black-box attacks. In particular, we want to observe whether these defenses designed against white-box attacks (viz. AutoAttack) and established benchmarks (viz. Robustbench), generalize to black-box attacks. A positive answer to these questions would maintain (and even raise) the motivation to develop defenses in white-box settings, whereas a negative answer would raise an urgent necessity of investigating specific defenses against black-box attacks.
4.1 Experimental Setup
We followed the same protocol as in the previous section, in terms of datasets, and adversarial black-box attacks. We selected nine robust target models from the Robustbench leaderboard. We ensured to have an even distribution across the robustness spectrum for the nine targets. The robust accuracy of the most robust model, a Swin-L (Liu et al., 2024) is 59.64% (that is, a success rate of AutoAttack of 25.97%), and the robust accuracy of the least robust model, a ResNet18 (Salman et al., 2020) is 25.32% (for a success rate of AutoAttack of 50.53%). Table 1 presents the architecture and the number of parameters of the models. These nine models have been defended using different categories of mechanisms, listed in Table 2.
We kept the same surrogate ensemble for BASES and TREMBA, and we used the surrogate model ResNet50 with standard training for single-surrogate attacks. In addition to the selected black-box attacks, we incorporated the SGM attack due to its compatibility with the ResNet50 surrogate model.
Table 1: Summary of different models from the RobustBench leaderboard (Croce et al., 2020) and their configurations. Legend of the defenses in Table 2.
Rank Reference Architecture Robust Accuracy Defense Architecture Type Parameters
1 Liu et al. (2024) Swin-L (Liu et al., 2021) 59.56 % A Transformer 196.53M
2 Bai et al. (2024) ConvNeXtV2+Swin-L (Liu et al., 2022; 2021) 58.50 % A, B Transformer & Convolution 394.49M
3 Liu et al. (2024) ConvNeXt-L (Liu et al., 2022) 58.48 % A Convolution 197.77M
5 Liu et al. (2024) Swin-B (Liu et al., 2021) 56.16 % A Transformer 87.77M
7 Liu et al. (2024) ConvNeXt-B (Liu et al., 2022) 55.82 % A Convolution 88.59M
12 Singh et al. (2024) ViT-S+ConvStem (Dosovitskiy et al., 2020; Xiao et al., 2021) 48.08 % C Transformer 22.78M
17 Salman et al. (2020) WideResNet50.2 (Zagoruyko, 2016) 38.14 % D Convolution 68.88M
18 Salman et al. (2020) Resnet50 (He et al., 2016) 34.96 % D Convolution 25.56M
21 Salman et al. (2020) Resnet18 (He et al., 2016) 25.32 % D Convolution 11.69M
Table 2: List of the defense mechanisms.
Label Description
A Adv training with large data augmentation, regularisation, weights averaging, pretraining
B Non linear ensemble of two base models robust and vanilla
C Downsample convolutional layers before subsequent network layers
D Standard adversarial training
4.2 Results
Figure 2: Success rate of blackbox attacks against SoTA defenses.
We analyze the performance of defenses against black-box attacks by examining the impact of different defense strategies, size of models, and architecture of models, as well as the relationship between AutoAttack success rates and black-box attacks success rates.
Impact of defense mechanism.
We categorize the defended models from Table 1 into four distinct families (A, B, C, and D), reflecting the various training strategies employed to enhance model robustness. In Figure 2 we show the ASR of the 12 black-box attacks against the four defenses (robust models). The models are sorted according to the Robustbench leaderboard, with stronger defenses against AutoAttack positioned to the right. Models using the same category of defense mechanisms but differing architectures are grouped in the same color. We observe that stronger defenses against AutoAttack lead to stronger defenses against black-box attacks. For instance, Defense A (red) outperforms Defense C (yellow), which in turn surpasses Defense D (blue) across all black-box attacks. An exception to this trend is observed with the ensemble Conv+Swin Defense, ranked #2 on Robustbench and mixing mechanisms of types A and B. Although ranked second against AutoAttack, it performs similarly to Defense D when subjected to black-box attacks. This is explained by the composition of defense mechanisms A and B, which combines the adversarially trained model from Defense A with a standard (vanilla) model, as explained in Section 2.2. This hybrid nature introduces a vulnerability: while this defense ranks among the best defenses against AutoAttack, the inclusion of a standard model component causes its performance against black-box attacks to align with that of weaker AutoAttack defenses. This vulnerability arises because black-box attacks often exploit vanilla surrogates, making the defense susceptible due to the standard model’s presence.
Table 3: Success rate on Resnet models trained by Salman et al. (2020).
Attacks Salman ResNet18 Salman ResNet50 Salman WideResNet50.2
AUTOATTACK 50.53 43.84 42.87
MI-FGSM 1.28 ± 0.00 1.68 ± 0.00 1.34 ± 0.00
DI-FGSM 1.41 ± 0.02 2.20 ± 0.10 1.99 ± 0.04
TI-FGSM 2.26 ± 0.04 2.81 ± 0.07 2.78 ± 0.10
VMI-FGSM 1.24 ± 0.05 1.89 ± 0.04 1.48 ± 0.05
VNI-FGSM 1.35 ± 0.02 1.72 ± 0.03 1.59 ± 0.05
ADMIX 1.68 ± 0.06 2.68 ± 0.13 2.12 ± 0.08
UAP 0.86 ± 0.10 1.02 ± 0.01 0.93 ± 0.07
SGM 1.25 ± 0.00 1.62 ± 0.00 1.43 ± 0.00
GHOST 1.21 ± 0.18 1.69 ± 0.05 1.49 ± 0.06
LGV 1.63 ± 0.13 1.72 ± 0.26 1.52 ± 0.08
BASES 2.36 ± 0.03 2.83 ± 0.04 2.35 ± 0.06
TREMBA 4.06 ± 0.06 3.56 ± 0.06 3.22 ± 0.06
Impact of model size.
We compare in Table 3 the success rates of the black-box attacks against similar adversarially trained models. These models do not benefit from advanced defense mechanisms (e.g., synthetic data augmentation, …) and only differ in size. While larger models with adversarial training tend to achieve higher robustness to AutoAttack, the increase in size does not lead to lower success rates. Similar insights are revealed in Table 4, where we compare the robustness of two sizes of ConvNext models and SWIN Transformers. Larger models do not consistently demonstrate stronger robustness to adversarial perturbations in black-box settings, although increased size improves robustness against AutoAttack.
Impact of model architecture.
Liu et al. (2024) demonstrate that the best robustness against AutoAttack can be achieved using indiscriminately Convnets or Transformers models. Our results in Table 4 confirm that for a fixed size, both architectures demonstrated similar robustness against black-box attacks.
Table 4: Success rate on very large models trained by Liu et al. (2024).
Attacks Liu ConvNeXt-B Liu Swin-B Liu ConvNeXt-L Liu Swin-L
AUTOATTACK 28.01 27.14 25.95 25.98
MI-FGSM 0.63 ± 0.00 0.37 ± 0.00 0.49 ± 0.00 0.56 ± 0.00
DI-FGSM 0.90 ± 0.02 0.63 ± 0.03 0.74 ± 0.09 0.85 ± 0.04
TI-FGSM 0.97 ± 0.02 0.79 ± 0.07 1.07 ± 0.03 1.14 ± 0.08
VMI-FGSM 0.74 ± 0.02 0.41 ± 0.01 0.58 ± 0.03 0.55 ± 0.01
VNI-FGSM 0.74 ± 0.04 0.45 ± 0.02 0.49 ± 0.02 0.64 ± 0.04
ADMIX 1.04 ± 0.04 0.72 ± 0.02 0.74 ± 0.01 0.87 ± 0.00
UAP 0.19 ± 0.04 0.16 ± 0.02 0.18 ± 0.06 0.21 ± 0.04
SGM 0.69 ± 0.00 0.37 ± 0.00 0.49 ± 0.00 0.56 ± 0.00
GHOST 0.63 ± 0.03 0.46 ± 0.07 0.49 ± 0.09 0.49 ± 0.04
LGV 0.57 ± 0.04 0.36 ± 0.05 0.45 ± 0.05 0.46 ± 0.04
BASES 1.38 ± 0.06 0.96 ± 0.02 1.16 ± 0.00 1.26 ± 0.02
TREMBA 1.34 ± 0.04 1.14 ± 0.03 1.47 ± 0.01 1.25 ± 0.01
Relationship between AutoAttack success rate and black-box success rate. As shown in Figure 3, the scatter plot illustrates the relationship between the ASR of AutoAttack and that of 12 black-box attacks. Each data point represents the ASR on the y-axis of a given black-box attack (indicated by a color) against a specific robust model (determined by its AutoAttack success rate on the x-axis).
Although the initial ASR of black-box attacks is low (less than 5%), we observe a clear downward trend as robust models improve their effectiveness against AutoAttack, until the best four robust models (with AutoAttack success rate ranging from 25.48% to 27.14%). This vulnerability stems from the lack of correlation between model size and robustness, as these models all employ defense (A) with diverse sizes and architectures, as shown in Table 1.
Furthermore, the best robust model (25.48% AutoAttack success rate) underperforms compared to the other top robust models against black-box attacks. This occurs because it uses the ensemble defense (B), which, as discussed earlier, combines a robust model with a standard model, making it vulnerable to black-box attacks that often exploit surrogates resembling standard models.
Figure 3: Relation between success rate of AutoAttack and the success rate of black-box attacks.
Insight 2 SoTA defenses against AutoAttack generally correlate with improved black-box robustness. However, having a large model size does not always guarantee better robustness. Additionally, an ensemble defense that achieves similar robustness to a single defense against AutoAttack may be less effective against black-box attacks, especially those using surrogate models similar to the ensemble’s components.
5 Robust surrogates to Improve Black-Box Attacks
The results from the previous section demonstrated that models robust against AutoAttack remain relevant (to some extent) against black-box attacks. A critical concern then is whether these robust "defenses" could not be used by the attacker to improve its black-box attack. Indeed, most of the black-box require designing a surrogate and can therefore leverage the recent advances in model defense to strengthen their attacks. We hypothesize that these new defenses raise security risks by providing stronger surrogates to attackers.
5.1 Experimental Setup
We adhered to the same protocol as in previous sections, maintaining consistency in datasets and adversarial black-box attacks. We considered various configurations regarding the robustness of both the surrogate and target models. For the target models, we selected highly robust models (Swin-L, ConvNeXtV2+Swin-L, ConvNext-L from (Liu et al., 2024) and (Bai et al., 2024)), a moderately robust model (ResNet18 from (Salman et al., 2020)), and a non-robust model (ResNet50). This way, we can measure the effects of using robust surrogates on targets with top defense (A), a simple defense (D) and no defense.
As for surrogate, we used robust and non-robust models. The selection of robust surrogates was based on choosing the highly ranked robust models that are not used as target and are compatible with the attack methods. Hence, we used the robust surrogate RaWideResNet-101-2 from (Peng et al., 2023) for single surrogate attacks, because it is the best defense in Robustbench that supports skip connections (which are essential for the GHOST attack). For ensemble surrogate attacks, we also added the robust surrogate Swin-B from (Liu et al., 2024), as it is the best robust model available, excluding the target models. For single, non-robust surrogate attacks, we used a vanilla WideResNet-101-2 in order to have the same architecture as in the robust surrogate, ensuring fair comparison. For BASES and TREMBA (who use an ensemble of surrogates), we kept the same ensemble of vanilla models as in our previous experiments. We excluded the SGM attack that does not support the WideResNet-101-2 surrogates used in our setup.
5.2 Results
We present in Figure 4 the success rate of black-box attacks using non-robust surrogates and robust surrogates against a non robust target and robust targets.
Figure 4: Success rates of black-box attacks using vanilla and robust surrogates against a vanilla target and robust targets.
Non-robust target: The results clearly show that using robust surrogate models is detrimental to attack success rates. Across all black-box attacks, non-robust surrogates consistently outperform robust surrogates, with an average improvement of 35.39 percentage points in attack success rates on the vanilla ResNet50 model. For instance, GHOST improves from 17.95% to 49.36%, BASES from 17.94% to 81.08%, and TREMBA from 67.54% to 89.56%. Even UAP that had the lowest success rate improves from 0.73% to 7.11%.
Robust target: By contrast, robust surrogates yield significant improvement in the success rate of black-box attacks when targeting robust models. Notably, attacks using robust surrogates on top-tier robust models (e.g., Liu ConvNeXt-L and Liu Swin-L) yield up to 15 times higher success rates than their non-robust counterparts. For instance, when using the GHOST attack to target the highly robust Swin-L model, the robust surrogate achieves a success rate of 8.16%, compared to just 0.53% with the non-robust surrogate. Similarly, BASES improves its success rate from 1.26% to 12.55% against the Swin-L model. The significant improvement of success rate is consistent across all defense mechanisms, and all attacks, except LGV, and UAP. We hypothesize that the initial training phase of the LGV models, where a high learning rate is used over additional epochs, potentially weakens the model’s robustness during the early sampling phase, which contrasts with GHOST networks that maintain stronger robustness after their creation. This hypothesis is explored in Appendix D.
The results highlight that the choice of surrogate model plays a critical role in the effectiveness of transfer-based black-box attacks, particularly in restricted environments where the robustness of the target is unknown. Robust models, by design, tend to secure blind spot regions in the input space in which vanilla models often fail to predict adversarial examples correctly. As a result, transfer attacks on robust surrogates aim to fool other regions of the input space that are not yet secured by adversarial training, making them better suited for attacking robust models. On the other hand, transfer attacks on vanilla surrogates focuses more on blind spot regions of the input space that are also likely to be mispredicted by other vanilla target models.
Insight 3 The choice of the surrogate model, robust or vanilla, impacts the success of black-box attacks. Robust surrogates are more effective against robust targets, while vanilla surrogates excel against vanilla targets. Attackers can therefore improve attack effectiveness by adaptively selecting surrogates based on the perceived robustness of the target model.
6 Conclusion and Perspectives
This research investigated the effectiveness of black-box adversarial attacks against SoTA and standard robust models using ImageNet, structured around three series of experiments, uncovering key insights.. Our first findings demonstrate that even advanced black-box attacks often struggle against a relatively simple adversarial defense. This suggests that adversarial training can be a highly effective baseline defense for real-world systems accessible via APIs. Consequently, understanding the limitations of current black-box attacks against robust models is crucial for developing more effective attack strategies. Based on our second experiments, we found that defense mechanisms optimized against AutoAttack often generalize effectively to resist black-box attacks, even though performance differences are less pronounced in black-box settings. This finding supports the relevance of AutoAttack as a benchmark for black-box robustness and suggests that current efforts focused on white-box defenses provide a meaningful degree of robustness in black-box scenarios as well – though ensemble defenses may not always translate their AutoAttack robustness to black-box robustness, especially against attacks using similar surrogate models as the ensemble. Finally, we observed that selecting a surrogate with a level of robustness aligned with that of the target can significantly impact the success rate of transfer-based attacks. This finding reveal the possibility for strategic selection of surrogate models when planning black-box attacks. Future work should consider designing black-box attacks that specifically target SOTA robust models to better reflect real-world defensive capabilities. Additionally, further exploration of why SoTA defenses optimized for AutoAttack are also relevant for black-box robustness could yield valuable insights into cross-attack resilience mechanisms. Expanding the study of transfer-based attacks to consider varying levels of robustness between surrogate and target models would also help to deepen our understanding of transferability dynamics. Finally, while this study focused on image classification tasks, future research could extend the analysis to other tasks, such as object detection or segmentation, to understand how black-box robustness may vary across applications. In conclusion, this study demonstrates that SoTA defenses are notably resilient against black-box attacks, underscoring the importance of developing more targeted attack strategies to effectively challenge these modern robust models. As the field advances, these insights can guide both the creation of more sophisticated black-box attacks and the design of improved defense mechanisms that support the deployment of robust AI systems across a wide range of applications.
Acknowledgment
This research is supported by the Luxembourg National Research Fund under the AFR grant. We extend our gratitude to Martin Gubri (ParameterLabs) for his invaluable support with the LGV implementation.
Appendix
Appendix A Hyperparameters of Transfer Attacks
Tables 5 and 6 provide a comprehensive overview of the hyperparameters used for the ten transfer-based adversarial attacks in our experiments. They categorize the key hyperparameters into two groups: common hyperparameters shared across most attacks and attack-specific hyperparameters that define the unique characteristics of each method. Most of the attack hyperparameters were kept at their default values, as specified in the repositories or libraries that provide the implementations of adversarial attacks (Kim, 2020; Woo, 2024).
A.1 Common Hyperparameters
Table 5 provides the default values and explanations for these common hyperparameters, such as the step size (Alpha), momentum factor (Decay), and number of iterations (Steps), which are essential for iterative attacks as they control the perturbation magnitude and the optimization process.
Table 5: Common hyperparameters for transfer-based attacks.
PARAMETER DEFAULT VALUE EXPLANATION APPLIES TO
Alpha 2/255 Step size or update rate for perturbation. All attacks except UAP
Decay 1.0 Momentum factor. All attacks except UAP
Steps 10 Number of iterations. All attacks except UAP
Epsilon (eps) 4/255 Maximum perturbation for adversarial example. All attacks
A.2 Attack-Specific Hyperparameters
In addition to the shared hyperparameters, each attack incorporates its own unique hyperparameters tailored to its specific mechanisms. Table 6 summarizes these attack-specific hyperparameters, listing the default values, explanations, and the corresponding attacks to which they apply.
Table 6: Attack-specific hyperparameters for transfer-based attacks.
PARAMETER DEFAULT VALUE EXPLANATION APPLIES TO
Resize Rate 0.9 Rate at which the input images are resized. DIFGSM, TIFGSM
Diversity Prob 0.5 Probability to apply input diversity. DIFGSM, TIFGSM
Kernel Name Gaussian Type of kernel used. TIFGSM
Kernel Length 15 Length of the kernel. TIFGSM
Sigma (nsig) 3 Radius of the Gaussian kernel. TIFGSM
N 5 Number of samples in the neighborhood. VMI, VNI
Beta 1.5 Upper bound for neighborhood. VMI, VNI
Prior Type no_data Type of prior used for attack optimization. UAP
Patience Interval 5 Number of iterations to wait before checking convergence. UAP
Gamma 0.5 Decay factor for gradients from skip connections. SGM
LGV Epochs 5 Number of epochs for training the LGV models. LGV
LGV Models Epoch 2 Number of models used per epoch. LGV
LGV Learning Rate 0.05 Learning rate for the LGV training phase. LGV
Randomized Modulating Scalar 0.22 Drawn from the uniform distribution 1-scalar, 1+scalar. GHOST
Portion 0.2 Portion for the mixed image. Admix
Size 3 Number of randomly sampled images. Admix
Appendix B Hyperparameters of Query Attacks
Tables 7 and 8 provides an overview of the hyperparameters used in BASES and TREMBA, respectively. Most of these hyperparameters were kept at their default values, as specified in their repositories, since our experiments aimed to evaluate their performance under standard configurations. For BASES, key hyperparameters include the ensemble size (N_wb), the type of fusion for surrogate models (Fuse), and the number of weight updates (Iterw). Similarly, TREMBA, which operates in two phases (training and attacking), relies on hyperparameters such as the number of training epochs (Epochs), the number of attack iterations (Num_Iters), the number of samples for Natural Evolution Strategy (NES), and latent space parameters like the Gaussian noise standard deviation (Sigma).
Table 7: Hyperparameters for the BASES attack.
PARAMETER DEFAULT VALUE EXPLANATION
Epsilon (eps) 4 Perturbation bound magnitude (0 to 255 pixel range)).
N_wb 10 Number of models in the ensemble.
Bound linf Perturbation bound norm type.
Iters 10 Number of inner iterations.
Fuse loss Fuse method, e.g., loss or logit.
Loss_name CW Loss function used for optimization.
Algo PGD White-box algorithm used for the perturbation machine
X 3 Factor to scale the step size alpha.
Learning Rate (lr) 0.005 Learning rate for weight updates.
Iterw 20 Number of iterations to update weights.
Table 8: Hyperparameters for the TREMBA attack.
PARAMETER DEFAULT VALUE EXPLANATION
Epsilon (eps) 8/255 (Training Phase), 4/255 (Attacking Phase) Perturbation bound magnitude.
Learning Rate (G) 0.01 (Training Phase), 5.0 (Attacking Phase) for gradient updates, for latent updates
Momentum 0.9 (Training Phase), 0.0 (Attacking Phase) for SGD, for attack on the embedding space.
Num_images 49k (Training Phase) Size of the training set
Epochs 500 Number of epochs for Training.
Schedule 10 Epochs between Training learning rate decay.
Gamma 0.5 Training Learning rate decay factor.
Margin 200.0 (Training Phase), 5.0 (Attacking Phase) Margin for the loss function.
Sample_Size 20 Number of samples for NES
Num_Iters 200 Number of iterations for NES.
Sigma 1.0 Standard deviation of the Gaussian noise for NES.
Learning Rate Min 0.1 Minimum learning rate in the Attacking Phase.
Learning Rate Decay 2.0 Learning rate decay factor in the Attacking Phase.
Plateau_Length 20 Number of iterations to monitor the objective function
Plateau_Overhead 0.3 Allowed loss threshold to change before the decay
Appendix C Increasing the Budget of Black-Box Attacks
We quadrupled (multiplied by 4) the budget of the best black-box attacks in section 3 (BASES and TREMBA) against the robust ResNet50 model, to reassure that black-box attacks reduced effectiveness (compared to their effectiveness against the vanilla ResNet50 model) is not due to a limited budget. For BASES, we increased the number of inner iterations of PGD within the perturbation machine (Iters) to 40. For TREMBA, we increased the Number of iterations for NES (Num_Iters) to 800.
Figure 5 presents the success rates of BASES and TREMBA attacks against the vanilla ResNet-50 model with standard attack budget, and against the adversarially trained ResNet-50 model under both standard and quadrupled attack budgets. Both attacks underperform against the adversarially trained ResNet50 under quadrupled budgets compared to the vanilla model with standard budgets. For instance, BASES achieves 2.56% ASR on the robust ResNet50 Budget (x4) versus 81.08% on the vanilla ResNet50. Similarly, TREMBA’s ASR is at 5.24% on the robust ResNet50 Budget (x4) compared to 89.56% on the vanilla ResNet50. These findings confirm that our earlier results in section 3 were not a consequence of budget limitations.
Figure 5: The blue bars show the success rate for the vanilla ResNet50 model, while the orange bars show the results for the robust ResNet50 model, and the green bars show the success rate for the ResNet50 model by quadrupling the budget of the attacks.
Appendix D AutoAttack Robustness of LGV and GHOST Models
To evaluate the impact of the LGV weight collection phase, and the GHOST perturbing phase on model robustness, we collected 10 models after finishing the first phase for both LGV and GHOST, prior to proceeding to the attacking phase. Specifically, the LGV variations were obtained by training the surrogate model for additional epochs using a high learning rate, during which the weights of the models were collected periodically. For GHOST networks, the variations were obtained by applying random perturbations to the skip connections of the surrogate model. Using the robust surrogate model WideResNet-101-2 introduced in (Peng et al., 2023) and described in Section 5, we evaluated the robustness of the resulting LGV models and compared it with the robustness of GHOST models. The robust accuracies for the collected models are reported in Table 9.
Given that both LGV and GHOST models are derived from the same robust surrogate model, we observed a difference in their robustness. The robust accuracies for all LGV models is 0% (indicating 100% attack success rates), while the robust accuracies for GHOST networks is on average 51.54% across the 10 collected models. This confirms that the additional training phase of the LGV models with a high learning rate decreases the robustness of the surrogate models. Consequently, this explains why the performance improvement for LGV models when using a robust surrogate against robust targets is lower compared to the GHOST networks. As a good practice, it is essential to evaluate the robustness of surrogate models prior to the attacking phase. This ensures that any preliminary training phase does not inadvertently weaken the surrogate robustness.
Table 9: Robust accuracies of LGV and GHOST models before the attacking phase.
Num 1 Num 2 Num 3 Num 4 Num 5 Num 6 Num 7 Num 8 Num 9 Num 10 Average
GHOST 51.39 51.86 51.74 51.02 51.58 50.76 52.39 52.81 51.11 50.78 51.54
LGV 0 0 0 0 0 0 0 0 0 0 0
Appendix E Detailed Tabular Results
We provide the tabular representation corresponding to the results presented in the main paper. Tables 10, 11, 12, 13 represent the success rates for section 3, 4, and 5 with vanilla surrogates and robust surrogates, respectively. These tables display the mean and standard deviation of success rates computed across three random seeds.
Table 10: Success rates of black-box attacks against one vanilla, and one robust ResNet50 model.
MODELS MI-FGSM DI-FGSM TI-FGSM VMI-FGSM VNI-FGSM ADMIX UAP GHOST LGV BASES TREMBA
Vanilla ResNet50 57.77 $\pm$0.00 71.31 $\pm$0.32 43.59 $\pm$ 0.17 66.79 $\pm$ 0.07 68.00 $\pm$ 0.18 99.95 $\pm$ 0.00 8.89 $\pm$ 0.20 64.53 $\pm$ 0.13 19.81 $\pm$ 1.46 81.08 $\pm$0.49 89.56 $\pm$0.09
Robust ResNet50 1.53 $\pm$ 0.00 1.94 $\pm$ 0.03 2.3 $\pm$ 0.01 1.66 $\pm$ 0.03 1.85 $\pm$ 0.04 2.54 $\pm$ 0.09 0.98 $\pm$ 0.10 1.68 $\pm$0.03 1.61 $\pm$0.10 2.83 $\pm$0.04 3.56 $\pm$0.06
Table 11: Success rates of white-box AutoAttack and black-box attacks against robust targets from Robustbench.
RANK MODELS AUTOATTACK MI-FGSM DI-FGSM TI-FGSM VMI-FGSM VNI-FGSM ADMIX UAP SGM GHOST LGV BASES TREMBA
21 Salman ResNet18 50.53 1.28 $\pm$0.00 1.41 $\pm$0.02 2.26 $\pm$ 0.04 1.24 $\pm$ 0.05 1.35 $\pm$ 0.02 1.68 $\pm$ 0.06 0.86 $\pm$ 0.10 1.25 $\pm$ 0.00 1.21 $\pm$0.18 1.63 $\pm$0.13 2.36 $\pm$0.04 4.06 $\pm$0.06
18 Salman ResNet50 43.84 1.68 $\pm$ 0.00 2.2 $\pm$ 0.10 2.81 $\pm$ 0.07 1.89 $\pm$ 0.04 1.72 $\pm$ 0.03 2.68 $\pm$ 0.13 1.02 $\pm$ 0.01 1.62 $\pm$ 0.00 1.69 $\pm$0.05 1.72 $\pm$0.26 2.83 $\pm$0.04 3.56 $\pm$0.06
17 Salman WideResNet50.2 42.87 1.34 $\pm$ 0.00 1.99 $\pm$ 0.04 2.78 $\pm$ 0.10 1.48 $\pm$ 0.05 1.59 $\pm$ 0.05 2.12 $\pm$ 0.08 0.93 $\pm$ 0.07 1.43 $\pm$ 0.00 1.49 $\pm$0.06 1.52 $\pm$0.08 2.35 $\pm$0.06 3.22 $\pm$0.06
12 Sing ViT-S, ConvStem 34.76 0.89 $\pm$ 0.00 1.41 $\pm$ 0.05 1.69 $\pm$ 0.05 0.97 $\pm$ 0.04 1.00 $\pm$ 0.00 1.51 $\pm$ 0.05 0.32 $\pm$ 0.06 0.95 $\pm$ 0.00 0.98 $\pm$0.03 0.72 $\pm$0.12 1.55 $\pm$0.00 1.54 $\pm$0.03
7 Liu ConvNeXt-B 28.01 0.63 $\pm$ 0.00 0.9 $\pm$ 0.02 0.97 $\pm$ 0.02 0.74 $\pm$ 0.02 0.74 $\pm$ 0.04 1.04 $\pm$ 0.04 0.19 $\pm$ 0.04 0.69 $\pm$ 0.00 0.63 $\pm$0.03 0.57 $\pm$0.04 1.38 $\pm$0.06 1.34 $\pm$0.04
5 Liu Swin-B 27.14 0.37 $\pm$ 0.00 0.63 $\pm$ 0.03 0.79 $\pm$ 0.07 0.41 $\pm$ 0.01 0.45 $\pm$ 0.02 0.72 $\pm$ 0.02 0.16 $\pm$ 0.02 0.37 $\pm$ 0.00 0.46 $\pm$0.07 0.36 $\pm$0.05 0.96 $\pm$0.02 1.14 $\pm$0.03
3 Liu ConvNeXt-L 25.95 0.49 $\pm$ 0.00 0.74 $\pm$ 0.09 1.07 $\pm$ 0.03 0.58 $\pm$ 0.03 0.49 $\pm$ 0.02 0.74 $\pm$ 0.01 0.18 $\pm$ 0.06 0.49 $\pm$ 0.00 0.49 $\pm$0.09 0.45 $\pm$0.05 1.16 $\pm$0.00 1.47 $\pm$0.01
2 Bai ConvNeXtV2-L, Swin-L 25.48 2.03 $\pm$ 0.00 2.65 $\pm$ 0.03 1.76 $\pm$ 0.05 2.40 $\pm$ 0.05 2.48 $\pm$ 0.03 2.55 $\pm$ 0.05 0.58 $\pm$ 0.05 2.00 $\pm$ 0.00 1.78 $\pm$0.16 1.01 $\pm$0.14 4.77 $\pm$0.13 2.94 $\pm$0.12
1 Liu Swin-L 25.98 0.56 $\pm$ 0.00 0.85 $\pm$ 0.04 1.14 $\pm$ 0.08 0.55 $\pm$ 0.01 0.64 $\pm$ 0.04 0.87 $\pm$ 0.00 0.21 $\pm$ 0.04 0.56 $\pm$ 0.00 0.49 $\pm$0.04 0.46 $\pm$0.04 1.26 $\pm$0.02 1.25 $\pm$0.01
Table 12: Success rates of black-box attacks using vanilla surrogates against a vanilla target and robust targets.
Models GHOST LGV BASES TREMBA VMI-FGSM TI-FGSM ADMIX VNI-FGSM DI-FGSM UAP MI-FGSM
Standard ResNet50 49.36 $\pm$ 0.14 15.98 $\pm$ 1.13 81.08 $\pm$ 0.49 89.56 $\pm$ 0.09 61.22 $\pm$ 0.18 41.5 $\pm$ 0.35 64.75 $\pm$ 0.17 61.10 $\pm$ 0.13 65.95 $\pm$ 0.75 7.11 $\pm$ 0.25 51.62 $\pm$ 0.0
Salman ResNet18 1.13 $\pm$ 0.08 1.37 $\pm$ 0.12 2.36 $\pm$ 0.04 4.06 $\pm$ 0.06 1.29 $\pm$ 0.05 2.04 $\pm$ 0.06 1.88 $\pm$ 0.08 1.35 $\pm$ 0.06 1.51 $\pm$ 0.03 0.83 $\pm$ 0.03 0.98 $\pm$ 0.0
Liu ConvNeXt-L 0.48 $\pm$ 0.09 0.45 $\pm$ 0.01 1.16 $\pm$ 0.0 1.47 $\pm$ 0.01 0.7 $\pm$ 0.02 1.19 $\pm$ 0.1 0.99 $\pm$ 0.03 0.8 $\pm$ 0.04 0.92 $\pm$ 0.05 0.18 $\pm$ 0.04 0.6 $\pm$ 0.0
Liu ConvNeXtV2-L+Swin-L 1.7 $\pm$ 0.19 1.0 $\pm$ 0.03 4.77 $\pm$ 0.13 2.94 $\pm$ 0.12 2.51 $\pm$ 0.02 2.08 $\pm$ 0.04 2.52 $\pm$ 0.06 2.45 $\pm$ 0.05 2.79 $\pm$ 0.07 0.63 $\pm$ 0.03 2.0 $\pm$ 0.0
Liu Swin-L 0.53 $\pm$ 0.03 0.45 $\pm$ 0.01 1.26 $\pm$ 0.02 1.25 $\pm$ 0.01 0.74 $\pm$ 0.02 1.22 $\pm$ 0.05 0.93 $\pm$ 0.05 0.74 $\pm$ 0.04 0.88 $\pm$ 0.03 0.18 $\pm$ 0.0 0.51 $\pm$ 0.0
Table 13: Success rates of black-box attacks using robust surrogates against a vanilla target and robust targets.
Models GHOST LGV BASES TREMBA VMI-FGSM TI-FGSM ADMIX VNI-FGSM DI-FGSM UAP MI-FGSM
Standard ResNet50 17.95 $\pm$ 0.36 7.79 $\pm$ 1.58 17.94 $\pm$ 0.00 67.54 $\pm$ 0.13 16.47 $\pm$ 0.02 8.48 $\pm$ 0.18 15.22 $\pm$ 0.11 16.51 $\pm$ 0.04 15.28 $\pm$ 0.21 0.73 $\pm$ 0.0 15.96 $\pm$ 0.0
Salman ResNet18 14.15 $\pm$ 0.22 2.04 $\pm$ 0.19 11.95 $\pm$ 0.00 10.58 $\pm$ 0.14 12.54 $\pm$ 0.02 12.39 $\pm$ 0.04 14.4 $\pm$ 0.17 12.34 $\pm$ 0.05 13.83 $\pm$ 0.14 0.79 $\pm$ 0.0 12.44 $\pm$ 0.0
Liu ConvNeXt-L 9.23 $\pm$ 0.13 0.42 $\pm$ 0.12 12.02 $\pm$ 0.00 4.94 $\pm$ 0.06 9.71 $\pm$ 0.01 7.28 $\pm$ 0.04 9.1 $\pm$ 0.06 9.46 $\pm$ 0.00 9.74 $\pm$ 0.0 0.36 $\pm$ 0.0 9.67 $\pm$ 0.0
Liu ConvNeXtV2-L+Swin-L 6.69 $\pm$ 0.08 0.88 $\pm$ 0.15 11.77 $\pm$ 0.01 5.33 $\pm$ 0.11 7.54 $\pm$ 0.03 5.09 $\pm$ 0.04 7.0 $\pm$ 0.02 7.57 $\pm$ 0.00 7.07 $\pm$ 0.09 0.37 $\pm$ 0.0 7.59 $\pm$ 0.0
Liu Swin-L 8.16 $\pm$ 0.09 0.55 $\pm$ 0.13 12.55 $\pm$ 0.00 4.4 $\pm$ 0.08 8.95 $\pm$ 0.01 6.78 $\pm$ 0.07 8.21 $\pm$ 0.01 8.89 $\pm$ 0.02 8.72 $\pm$ 0.04 0.31 $\pm$ 0.0 8.94 $\pm$ 0.0

### v2

@@
 Although adversarial robustness has been extensively studied in white-box settings, recent advances in black-box attacks (including transfer- and query-based approaches) are primarily benchmarked against weak defenses, leaving a significant gap in the evaluation of their effectiveness against more recent and moderate robust models (e.g., those featured in the Robustbench leaderboard).
-In this paper, we question this lack of attention from black-box attacks to robust models.
-We establish a framework to evaluate the effectiveness of recent black-box attacks against both top-performing and standard defense mechanisms, on the ImageNet dataset.
-Our empirical evaluation reveals the following key findings: (1) the most advanced black-box attacks struggle to succeed even against simple adversarially trained models; (2) robust models that are optimized to withstand strong white-box attacks, such as AutoAttack, also exhibits enhanced resilience against black-box attacks; and (3) robustness alignment between the surrogate models and the target model plays a key factor in the success rate of transfer-based attacks.
-1 Introduction
-Since the discovery that deep learning models are susceptible to minor input disturbances, yielding adversarial examples (AE) (Goodfellow et al., 2015), the development of robust models has become one of the most active topics in the machine learning community.
-This topic has been thoroughly explored for white-box settings, leading to standardized attacks (Croce & Hein, 2020b) and benchmarks (Croce et al., 2020).
-However, realistic evaluations of adversarial attacks remain less explored.
-It involves scenarios where the attacker has limited or no knowledge of the target model to be attacked.
-These scenarios are referred to as black-box (or gray-box with partial knowledge (Guo et al., 2018)).
-Black-box attacks belong to one of the two following settings: When the output of the model provides both category labels and scores, it is referred to as a score-based black-box attack.
-Conversely, if the output includes only category labels, it is named a decision-based black-box attack (Brendel et al., 2018).
-Given that most computer vision Application Programming Interface (APIs) (Google, 2023; Imagga, 2023) generally provide image prediction categories alongside scores, most research in a black-box setting considers score-based black-box attacks.
-Hence, this work focuses on this category.
-State of the art (SOTA) black-box attacks leverage one or both of the following mechanisms: (1) adversarial examples transferability and (2) iterative queries with meta-heuristics.
+In this work, we argue that this gap is problematic and the previous benchmarks conclusions do not hold under more robust evaluation frameworks, leading to contradicting conclusions on the transferability of adversarial attacks.
+We extensively evaluate the effectiveness of 13 popular black-box attacks, representative of the top ten popular transferability theories.
+We benchmark these attacks against eight top-performing and standard defense mechanisms on the ImageNet dataset.
+Our empirical evaluation reveals the following key findings: (1) the most advanced black-box attacks struggle to succeed even against simple adversarially trained models; (2) robust models that are optimized to withstand strong white-box attacks, such as AutoAttack, also exhibit enhanced resilience against black-box attacks; and (3) robustness alignment between the surrogate models and the target model can significantly impact the success rate of transfer-based attacks.
+Index Terms:
+adversarial attack, robustbench, defenses.
+[…]
+I Introduction
+TABLE I: Theories about the transferability of adversarial examples.
+In bold when the evaluation protocol of the black-box is sufficiently robust.
+We report the largest image dataset, the smallest perturbation distance, and the deep learning model families used in the evaluation.
+FNN for feed-forward neural network, CNN for convolutional neural network
+Theory Dataset Representative attacks Attack budgets Target architectures Target defenses
+Linearity [27] ImageNet FGSM $L_{2}=0.25$ Convolution FGSM AT
+Decision Boundary Similarity [44] ImageNet FGSM $L_{0}=23$ Convolution $\varnothing$
+Gradient Similarity [19] MNIST FGSM $L_{2}=1$ FNN $\varnothing$
+Adversarial Subspace [68] MNIST FGSM $L_{2}=4$ FNN, CNN FGSM AT
+Non-Robust Features [37] ImageNet PGD $L_{2}=3$ CNN PGD AT
+Model complexity [76] ImageNet IFGSM $L_{2}=15$ CNN $\varnothing$
+Perturbation interactions [71] ImageNet PGD $L_{\infty}=16/255$ CNN FGSM AT
+Spectral properties [14] CIFAR-10 PGD $\mathbf{L_{\infty}=2/255}$ FNN, CNN PGD AT
+Knowledge transferability [42] STL-10 PGD $\mathbf{L_{2}=2/255}$ CNN $\varnothing$
+Generalization gap [72] CIFAR-100 PGD $L_{\infty}=8/255$ CNN $\varnothing$
+Proposed protocol ImageNet AutoAttack, BASES $\mathbf{L_{\infty}=4/255}$ CNN, Transformers Robustbench
+Since the discovery that deep learning models are susceptible to minor input disturbances, resulting in adversarial examples (AE) [27], the development of robust models has become one of the most active topics in the machine learning community.
+This topic has been thoroughly explored for white-box settings, settings where attackers have full knowledge of the target, leading to powerful standardized attacks [18] and benchmarks [15].
+In parallel, more ”realistic” studies of adversarial vulnerability involved restricting the knowledge or the capabilities of the attacker.
+These scenarios are referred to as black-box scenarios (or gray-box when partial knowledge of the target is available [30]).
+State-of-the-art (SoTA) black-box attacks leverage one or both of the following mechanisms: (1) adversarial examples transferability and (2) iterative queries with meta-heuristics.
 Transferability methods craft adversarial examples on a specifically built surrogate model (with white-box access) such that this example should transfer to (i.e.
@@
 Iterative query methods use query feedback from the target model to guide or refine the optimization process to uncover its vulnerabilities.
-Transferability is difficult to achieve when the target differs from the surrogate (in terms of architecture, training, or dataset) and the factors behind transferability (or lack thereof) remain an active field of study (Charles et al., 2020; Dong et al., 2018a; Li et al., 2018; Wu et al., 2020a).
-This is because adversarial attacks crafted on the surrogate lead to examples that maximize the loss function of the surrogate model (Goodfellow et al., 2015; Kurakin et al., 2017), while the target model has a different loss function.
-Methods to improve transferability mostly rely on building diversity during the optimization of adversarial examples (Li et al., 2018; Wu et al., 2020a; Xie et al., 2019a).
-Iterative-query attacks also face multiple challenges: With increased search space (large images, for example), common query attacks show a significant decrease in the Attack Success Rate (ASR) or require large amounts of queries to achieve sufficient ASR (Mohaghegh Dolatabadi et al., 2020; Huang & Zhang, 2020; Feng et al., 2022).
-BASES (Cai et al., 2022b), a recent Black-box Surrogate Ensemble Search attack achieved more than 95% ASR with a few queries (<5) on average across a large scope of non-robust models.
-Though the initial motivation for black-box attacks and their evaluation protocol is to improve the realism of robustness evaluations, they generally suffer from a common pitfall: they do not consider robust models which have already demonstrated their improved performance against white-box attacks.
-We hypothesize that the evaluation results of black-box attacks can be misleading because even simple robustification mechanisms can be sufficient to successfully evade black-box attacks.
-Going further, we want to assess whether innovations in defenses provide positive benefits against black-box attacks like they did in white-box settings.
-To the best of our knowledge, this paper is the first to study the effectiveness of black-box attacks against strandardized defenses to demonstrate the need to confront black-box attacks to existing defenses.
+Black-box attacks are powerful tools to study real-world attack scenarios and to better understand the mechanisms behind adversarial vulnerabilities of deep learning models, in particular through the study of attacks’ transferability.
+However, they still rely on outdated and unstandardized evaluation practices (cf.
+Table I).
+In contrast to white-box research, recent black-box attacks are still designed and evaluated using large epsilon noise budgets, small architectures, and non-robust target models.
+Traditional theories such as the decision boundary similarity [44], the adversarial subspace theory [68], and the non-robust features [37] are still well recognized, but have been demonstrated using simple transfer attacks, with very large budgets and non-robust models.
+In contrast, Robustbench [15] recommended much stronger models and enforced a budget of $\epsilon=4/255$ as a reasonable evaluation budget.
+At least ten of the main theories have been empirically demonstrated using incomplete black-box attacks protocols (Table I): linearity [27], gradient similarity [19], adversarial subspace [68], non-robust features [37], model complexity [76], adversarial perturbation interactions [71], spectral properties [14], knowledge transferability [42], generalization gap [72], and decision boundary similarity [44, 34].
+Our work does not intend to challenge all previous work and theories about the vulnerability of deep learning models to adversarial attacks; indeed, many theories were evaluated with the best protocols at that time.
+We would rather like to demonstrate that adhering to more stringent evaluation protocols leads to different insights and conclusions on the vulnerability of the models and effectiveness of the black-box attacks.
+In addition, given the importance of black-box attacks in assessing real-world robustness of ML systems put in production, we also hypothesize that the previous evaluation results of black-box attacks can be misleading.
+Even old robustification mechanisms enable us to successfully evade supposedly SoTA black-box attacks.
+LGV [28], one of the strongest transfer attacks, sees its success rate drop from $87.39\%$ to $2.17\%$ under a 2020 defense by Salman et al.[59].
+Our first insight suggests to reconsider previous claims of effectiveness of black-box attacks and demonstrates that simple adversarial training is sufficient against the majority of SoTA black-box attacks.
+While our first insight challenges the effectiveness of SoTA attacks, we tackle the benefits of SoTA defenses in our second study.
+We evaluated 9 defenses from Robustbench white-box ImageNet benchmark [16] and categorized these defenses into four families.
+Our analysis demonstrates that within the same family of defenses, architectures and model size have limited impact on the effectiveness of defenses.
+Our main insight from this study is that SoTA defenses are starting to plateau and show diminishing returns against SoTA black-box attacks.
+Finally, our work argues that an incomplete evaluation of the interactions between new adversarial defenses and black-box attacks can actually lead new defenses to strengthen already existing black-box attacks, opening new surface attacks that were not considered before.
+For example, BASES [8] can leverage recent robust models from Liu et al.[43] to increase its effectiveness against the most robust models from $1.26\%$ to $12.55\%$.
+In fact, we analyzed the top 20 defenses in the Robustbench ImageNet leaderboard ($L_{\infty}$, eps=4/255), and none have explored how their design choices might inadvertently aid black-box attacks.
+Even Blackboxbench, the most recent and complete benchmark of black-box attacks [85], did not consider the interactions between SoTA white-box defenses and SoTA black-box attacks.
+Our last insight highlights the new risks posed by SoTA defenses, that can be leveraged to improve existing (weak) attacks.
+To the best of our knowledge, this paper is the first to study the effectiveness of black-box attacks against standardized protocols and to demonstrate the need to confront black-box attacks to robust defenses and up-to-date evaluation settings.
 Our contributions can be summarized as follows:
@@
   2. 2.
-    We show that white-box robustness could be served as a proxy for black-box robustness. Defenses optimized against AutoAttack generalize well to black-box scenarios.
+    We show that white-box robustness could serve as a proxy for black-box robustness. Defenses optimized against AutoAttack generalize well to black-box scenarios.
   3. 3.
     We demonstrate that these effective defense mechanisms can inadvertently contribute to enhancing the success rate of black-box attacks.
-By using robust models as surrogates, attacks can generate adversarial examples more likely to transfer to robust models, leading to an average increase in success rate of 6.42 percentage points across attacks and target models.
-Our contributions send a clear message to the adversarial research community regarding the importance of considering white-box defenses in the evaluation of black-box attacks.
-They also urge the researchers working on defense mechanisms to study how their defenses can help improve transfer-based attacks.
-2 Related Work
-2.1 Black-Box Attacks: Transfer-Based and Query-Based Methods
-Black-box attacks have seen notable advancements, especially within transfer-based and query-based approaches.
+By using robust models as surrogates, attacks can generate adversarial examples more likely to transfer to robust models, leading to an average increase in success rate of 6.49 percentage points across attacks and target models.
+Our contributions send a clear message to the adversarial research community on the importance of considering white-box defenses when designing or evaluating black-box attacks, or exploring transferability mechanisms and theories.
+They also urge researchers working on defense mechanisms to consider how their recent defenses can help malicious parties improve existing black-box attacks, and study how that their work could lead to stronger attacks against seemingly robust defenses.
+We provide a replication package with our experiments and benchmark on https://figshare.com/projects/RobustBlack/265786.
+II Background
+II-A Preliminaries
+Adversarial perturbation [15]
+Let $x\in\mathbb{R}^{d}$ be an input point and $y\in\{1,\ldots,C\}$ be its correct label.
+For a classifier $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{C}$, we define a successful adversarial perturbation with respect to the perturbation set $\Delta\subseteq\mathbb{R}^{d}$ as a vector $\vec{\delta}\in\mathbb{R}^{d}$ such that
+$\mathop{\rm arg\,max}\limits_{c\in\{1,\ldots,C\}}f({x}+\delta)_{c}\neq y\quad\textrm{and}\quad{\delta}\in\Delta,$
+where the perturbation set $\Delta$ is chosen such that all examples in $x+\delta$ have $y$ as their true label.
+This motivates a robustness measure called robust accuracy, which is the fraction of data points on which the classifier $f$ predicts the correct class for all possible perturbations from the set $\Delta$.
+Computing the exact Robust Accuracy (RA) is in general intractable and, when considering $\ell_{p}$-balls as $\Delta$, NP-hard even for single-layer neural networks.
+In practice, an upper bound for robust accuracy is determined by adversarial attacks, generally involving optimization of a differentiable loss function or a reward through search algorithms that aim to identify a successful adversarial perturbation.
+The tightness of the upper bound depends on the strength of the attack.
+Success rate
+Given that different models show varying clean performances (e.g., test accuracy on the original set), robust accuracy will be impacted by the initial performance as much as the intrinsic robustness of the models.
+Thus, we base our study on an agnostic test performance metric: attack success rate (ASR).
+We define ASR for a classifier $f$ under attack $\mathcal{A};\mathcal{A}(x)={x}+\delta$ as
+$\text{ASR}(f,\mathcal{A})=\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\mathbb{I}\left(f(\mathcal{A}(x))\neq f(x)\right)\right].$
+For a model with a 100% test accuracy, the two metrics are related $\text{ASR}=1-\text{RA}$.
+II-B Black-Box Attacks: Transfer-Based and Query-Based Methods
+Black-box attacks have seen notable advances, especially within transfer-based and query-based approaches.
 Transfer-based attacks rely on adversarial examples generated by a surrogate model, which are then tested on a target model.
-Early approaches like Projected gradient descent (PGD) (Madry et al., 2017) laid the groundwork for such attacks by iteratively optimizing perturbations to maximize the success rate on the target model .
-This foundational method was enhanced by the Momentum Iterative Fast Gradient Sign Method (MI-FGSM), which incorporated momentum into PGD to improve transferability and stability across iterations (Dong et al., 2018b).
-Further innovations, such as the Diverse Input Fast Gradient Sign Method (DI-FGSM) and Translation-Invariant FGSM (TI-FGSM), introduced input transformations and image translations, respectively, to increase the robustness and generalizability of adversarial examples across different models (Xie et al., 2019b; Dong et al., 2019).
-Subsequent methods like Variance Tuning (VMI, VNI) refined gradient calculations by considering gradient variance across iterations (Wang & He, 2021), leading to more effective adversarial perturbations.
-Similarly, ADMIX leveraged diverse inputs by mixing the target image with randomly sampled images to generate more transferable attacks (Wang et al., 2021).
-Skip Gradient Method (SGM) focused on model architecture, by using skip connections which facilitate the creation of adversarial examples that have high transferability (Wu et al., 2020b).
-Universal Adversarial Perturbations (UAP) took a different approach by focusing on universality, generating a single small perturbation that can effectively disrupt a classifier across most natural images (Moosavi-Dezfooli et al., 2017).
-Notably, Ghost networks (GHOST) and Large Geometric Vicinity (LGV) contributed to transferability by modifying skip connections in surrogate models and by exploring variations in surrogate models vicinity, respectively (Li et al., 2020; Gubri et al., 2022).
-In query-based attacks, which generally have higher success rates than transfer attacks, methods like Zeroth Order Optimization (ZOO), Decision Boundary Attack (DBA), Natural Evolutionary Strategies (NES),and HopSkipJump rely on iterative query feedback to optimize perturbations, often not utilizing surrogate models to boost efficiency (Brendel et al., 2018; Chen et al., 2017; Ilyas et al., 2018; Chen et al., 2020).
-Recent work, such as BASES and Transferable Model-based Embedding (TREMBA) (Cai et al., 2022a; Huang & Zhang, 2019), integrates both surrogate representations and query feedback to create highly effective black-box attacks.
+Early approaches like Projected Gradient Descent (PGD) [47] laid the groundwork for such attacks by iteratively optimizing perturbations to maximize the success rate on the target model.
+This foundational method was enhanced by the Momentum Iterative Fast Gradient Sign Method (MI-FGSM), which incorporated momentum into PGD to improve transferability and stability across iterations [22].
+Further innovations, such as the Diverse Input Fast Gradient Sign Method (DI-FGSM) and Translation-Invariant FGSM (TI-FGSM), introduced input transformations and image translations, respectively, to increase the robustness and generalizability of adversarial examples across different models [79, 23].
+Subsequent methods such as Variance Tuning (VMI, VNI) refined gradient calculations considering gradient variance between iterations [69], leading to more effective adversarial perturbations.
+Similarly, ADMIX leveraged diverse inputs by mixing the target image with randomly sampled images to generate more transferable attacks [70].
+The Skip Gradient Method (SGM) focused on model architecture, using skip connections which facilitate the creation of adversarial examples that have high transferability [75].
+Universal Adversarial Perturbations (UAP) took a different approach by focusing on universality, generating a single small perturbation that can effectively disrupt a classifier in most natural images [51].
+In particular, Ghost networks (GHOST) and Large Geometric Vicinity (LGV) contributed to transferability by modifying skip connections in surrogate models and by exploring variations in surrogate models’ vicinity, respectively [41, 28].
+In query-based attacks, which generally have higher success rates than transfer attacks [2], methods like Zeroth Order Optimization (ZOO) [6], Decision Boundary Attack (DBA) [36], Natural Evolutionary Strategies (NES) [9], HopSkipJump [9], SignFlip [12], and Square [3] rely on iterative query feedback to optimize perturbations, often not utilizing surrogate models to boost efficiency .
+The following attacks then focused on reducing the number of queries.
+[2] introduced a query-efficient black-box attack using randomized square-shaped updates at image boundaries, SimBA [29] proposed a simple attack using orthogonal search directions (e.g., DCT basis).
+Sign-OPT [13] introduces a hard-label attack estimating gradient signs instead of magnitudes, while SignHunt [1] uses sign bits for gradient estimation.
+to minimize the number of queries.
+RayS [10] reformulates the boundary search as discrete optimization, eliminating gradient estimation.
+Recent work combined both surrogate representations and query feedback to create highly effective and efficient black-box attacks.
+Representative attacks include Transferable Model-based Embedding (TREMBA) and BASES [33, 7].
 BASES leverages an ensemble of surrogates with query feedback to dynamically adjust surrogate weighting, while TREMBA utilizes a generator trained with surrogate models, exploring latent space for more effective query attacks.
-2.2 Defenses Against AutoAttack
+II-C Defenses Against Strong Adversarial Attacks
 In parallel, adversarial defenses have evolved to mitigate the risks posed by increasingly sophisticated white-box attacks.
 As defenses became more robust, the need for a reliable, standardized benchmark to accurately assess their effectiveness grew increasingly apparent.
-In response, AutoAttack emerged as a reliable, parameter-free method for evaluating adversarial robustness, offering a computationally affordable, and a standardized benchmark applicable to various models (Croce & Hein, 2020b).
-Building on this, Robustbench established a leaderboard specifically for evaluating adversarial robustness, as outlined by (Croce et al., 2020).
-This leaderboard includes results on the ImageNet large-scale datasets (Deng et al., 2009), and offers a ranking comparison of various defense strategies across specific perturbation budgets.
-Among the defenses featured on the leaderboard are those proposed by (Salman et al., 2020; Singh et al., 2024; Liu et al., 2024; Bai et al., 2024; Madry et al., 2017), offering a thorough assesment on the effectiveness of the leading defense mechanisms.
-Traditional defenses, such as Madry’s adversarial training, initially combined clean and adversarial data during training, establishing a baseline for robustness (Madry et al., 2017).
+In response, AutoAttack emerged as a reliable, parameter-free method to evaluate adversarial robustness, offering a computationally affordable and a standardized benchmark applicable to various models [18].
+Building on this, Robustbench established a leaderboard specifically to evaluate adversarial robustness, as outlined by [15].
+This leaderboard includes results on ImageNet large-scale datasets [20], and offers a ranking comparison of various defense strategies in specific perturbation budgets.
+Among the defenses featured on the leaderboard are those proposed by [59, 61, 43, 5, 47], which offers a thorough assessment of the effectiveness of the leading defense mechanisms.
+Traditional defenses, such as Madry’s adversarial training, initially combined clean and adversarial data during training, establishing a baseline for robustness [47].
 Recent advancements have incorporated complex pre-training schemes and data augmentations (e.g., ConvStem architectures, RandAugment, MixUp, and CutMix).
-Singh et al.
-(2024) made alterations to the convnets architecture, particularly substituting PatchStem with ConvStem, and modified the training scheme to enhance the robustness against previously unencountered l1 and l2 threat models.
-Liu et al.
-(2024) further advances the training scheme with large-scale pre-training, combined with label smoothing, weight decay, and Exponential Moving Averaging (EMA) to enhance generalization against unseen adversarial examples.
-Bai et al.
-(2024) builds on a giving defense framework (Liu et al., 2024), focusing specifically on balancing clean and robust accuracy through the nonlinear mixing of robust and standard models, addressing the often-seen tradeoff between clean and robust performance.
-2.3 Robustness Under Black-Box Setting
+[61] made alterations to the convnets architecture, particularly substituting PatchStem with ConvStem, and modified the training scheme to enhance the robustness against previously unencountered l1 and l2 threat models.
+[43] further advances the training scheme with large-scale pre-training, combined with label smoothing, weight decay, and Exponential Moving Averaging (EMA) to enhance generalization against unseen adversarial examples.
+[5] builds on a given defense framework [43], focusing specifically on balancing clean and robust accuracy through the nonlinear mixing of robust and standard models, addressing the often-seen tradeoff between clean and robust performance.
+II-D Robustness Under Black-Box Setting
 In this part, we focus on how our work introduces a novel contribution compared to previous studies on defenses against black-box adversarial attacks in image classification tasks.
-We discuss four representative works (Dong et al., 2020; Mahmood et al., 2021a; b; Ghaffari Laleh et al., 2022), highlighting their key contributions and how our approach builds upon or complements these efforts by addressing gaps and introducing new perspectives in this domain.
-Dong et al.
-(2020) evaluated adversarial robustness in image classification tasks against white-box and black-box attacks.
+We discuss the following representative works [53, 21, 48, 49, 26, 85], highlighting their key contributions and how our approach builds on or complements these efforts by addressing gaps and introducing new perspectives in this domain.
+[53] introduced a black-box attack strategy targeting deep neural network (DNN) models and evaluated its effectiveness against adversarial training and defensive distillation.
+They varied the magnitude of the perturbation used during training and the attack phase and demonstrated that small perturbations of adversarial training in training led to gradient masking, which their attack could bypass by doubling the perturbation budget in the attacking phase, whereas larger perturbations in training improved robustness.
+In contrast, our work focuses on evaluating adversarially trained SOTA defenses against black-box attacks using a fixed small perturbation 4 over 255 in the attack phase.
+[21] evaluated adversarial robustness in image classification tasks against white-box and black-box attacks.
 Their work addressed various defense techniques such as robust training, input transformation, randomization, certified defenses, and model ensembling.
@@
 Moreover, our work explores a broader landscape of black-box attacks by including recent methods such as ADMIX, BASES, and TREMBA.
-In doing so, we provide an analysis of leading robust training defenses against recent black-box adversarial strategies.
-Mahmood et al.
-(2021a) noted that the majority of existing defenses primarily address white-box attacks, neglecting the crucial aspect of black-box adversarial robustness.
-They provided a wide evaluation of adversarial defenses with a Convolutional neural network (CNN) architecture on benchmark datasets such as CIFAR-10 (Krizhevsky, 2009) and Fashion-MNIST (Xiao et al., 2017) against 12 attacks.
-In a follow-up work, Mahmood et al.
-(2021b) extended this investigation to Vision Transformers (ViTs), evaluating their adversarial robustness against white-box attacks and two black-box attacks on ImageNet.
+In doing so, we provide an analysis of the leading robust training defenses against recent black-box adversarial strategies.
+[48] noted that the majority of existing defenses primarily address white-box attacks, neglecting the crucial aspect of black-box adversarial robustness.
+They provided a wide evaluation of adversarial defenses with a Convolutional Neural Network (CNN) architecture on benchmark datasets such as CIFAR-10 [40] and Fashion-MNIST [77] against 12 attacks.
+In a follow-up work, [49] extended this investigation to Vision Transformers (ViTs), evaluating their adversarial robustness against white-box attacks and two black-box attacks on ImageNet.
 They suggested that ensemble defenses can enhance robustness when attackers lack access to model gradients.
@@
 We leverage SoTA robust models to challenge the attacks used to evaluate robustness.
-In the field of computational pathology, Ghaffari Laleh et al.
-(2022) examined adversarial robustness with a focus on CNN and transformer models.
-This work addressed certain adversarial attacks (including Fast Adaptive Boundary (FAB) (Croce & Hein, 2020a), and Square (Andriushchenko et al., 2020)), as well as adversarial training using (PGD) with Dual-Batch Normalization (DBN).
-Our work extends this line of inquiry by incorporating SOTA adversarially trained models and a more comprehensive suite of black-box attacks, providing additional context for these attacks by increasing their number and offering a more challenging confrontation against SOTA defenses.
-3 Success Rate of Black-Box Attacks Against Adversarial Training
-As preliminaries, we investigate the effectiveness of 12 black-box adversarial attacks against Madry adversarial training (Madry et al., 2017) on a single, relatively simple ResNet50 model (He et al., 2016).
+In the field of computational pathology, [26] examined adversarial robustness with a focus on CNN and transformer models.
+This work addressed certain adversarial attacks (including Fast Adaptive Boundary (FAB) [17], and Square [4]), as well as adversarial training using (PGD) with Dual-Batch Normalization (DBN).
+Our work extends this line of inquiry by incorporating adversarially trained SOTA models and a more comprehensive suite of black-box attacks, providing additional context for these attacks by increasing their number and offering a more challenging confrontation against SoTA defenses.
+Zheng et al.[85] presented the BlackboxBench benchmark, which offers a comprehensive study for evaluating black-box adversarial attacks.
+Specifically, they implemented 59 black-box attack methods, including 29 query-based and 30 transfer-based algorithms.
+They conducted evaluations on two datasets (CIFAR-10 and ImageNet) and examined four attack settings (targeted and untargeted attacks under $L_{2}$ and $L_{\infty}$ norms).
+They explored both regular and adversarially trained models .
+Our work expands the evaluation of target and surrogate models by assessing black-box attacks with the most recent SoTA adversarially trained models from the RobustBench leaderboard.
+We revisit the role of alignment between surrogate and target models, confirming that training-strategy compatibility enhances transferability, a trend also observed in BlackboxBench, where we updated this analysis with SoTA defenses both as surrogates and as targets.
+A recent line of research investigated preprocessing and randomness injection defenses against black-box attacks.
+[55] proposes adding Gaussian noise to queries to disrupt gradient estimation in black-box attacks, combining it with Gaussian-augmentation fine-tuning for improved robustness-accuracy trade-offs; [11] proposed the Adversarial Attack on Attackers (AAA) defense.
+It introduces output logit perturbation to misdirect score-based attacks while preserving clean accuracy and improving calibration.
+[52] showed that randomizing hidden features provides better robustness than input randomization against query-based attacks, with plug-and-play implementation and minimal accuracy loss.
+[62] demonstrated that unaware preprocessing reduces attack efficacy by 7×, and designed preprocessor-aware attacks that easily overcome such defenses.
+II-E Explaining Transferability and Robustness in Black-Box Setting
+Various research investigated the factors behind the transferability (or lack-off) of adversarial examples. In our study, we covered each theory with at least one representative method.
+Decision Boundary Similarity [44]
+Models trained on the same task develop similar decision boundaries.
+Adversarial perturbations are highly aligned with weight vectors across models.
+Attacks relying on these mechanisms include PGD and HopSkipJump.
+Gradient Similarity [19]
+Two main factors contribute to transferability: (1) similarity between gradient directions of source and target models, and (2) smoothness of the loss landscape.
+Higher gradient similarity and lower variance in loss landscapes lead to increased transferability.
+Representative attacks include MI-FGSM and Variance Tuning (VMI/VNI).
+Shared Adversarial Subspaces [68]
+Adversarial examples span contiguous subspaces of large dimensionality.
+When different models are trained on the same task, a fraction of their adversarial subspaces overlap.
+DI-FGSM, Ghost Networks, and LGV use these mechanisms.
+Non-Robust Features [37]
+Models learn both robust features (perceivable by humans) and non-robust features (imperceptible but statistically correlated with labels).
+Adversarial examples exploit non-robust features.
+ADMIX and NES leverage this theory.
+Model Complexity/Capacity [76]
+Model-specific factors, including architecture and capacity, influence transferability.
+Adversarial examples generated from simpler/shallower models tend to transfer better to complex models (e.g.
+SGM).
+Interaction-Based [71]
+There is a negative correlation between adversarial transferability and interactions inside adversarial perturbations.
+Less interaction leads to higher transferability.
+TI-FGSM falls into this category.
+Knowledge Transferability [42]
+Models with high knowledge transfer (e.g., via fine-tuning) exhibit stronger adversarial example transfer. The embeddings trained in TREMBA leverage this knowledge transfer.
+Flatness/Manifold [25]
+There are conflicting results that higher flatness of adversarial examples enables better cross-model transferability.
+BASES, through its search algorithm, finds optimal weights to explore the manifold.
+Other researchers specifically explored the link between white-box robustness and transferability.
+[63] investigated the link between the robustness of the source models and the robustness of the target models for transferable attacks.
+Their results show as expected that adversarial examples generated using non-robust networks do not transfer to the adversarially trained networks.
+However, what they consider as robust source models is a simple Resnet50 with vanilla adversarial training, which has since been shown to only be slightly robust [15].
+Our study, on the other hand, leverages state-of-the-art robustification mechanisms and explores a large set of architectures for robust surrogates and targets.
+Although our initial results match, our study highlights further insights (e.g., using robust sources is actually detrimental when targeting non-robust targets).
+[84] confirmed the impact of model smoothness and gradient similarity by exploring the impact of weakly adversarial training (that is, training with mildly perturbed examples).
+The approach used to train the model, including data augmentation, synthetic data, regularization are considered in some of the models of our study.
+Our work focuses on robustification mechanism using extreme augmentations.
+All in all, our results are complementary; as they explore ”mildly robust models” and we focus on ”extremely robust models”.
+III Success Rate of Black-Box Attacks Against Adversarial Training
+As preliminaries, we investigate the effectiveness of 13 black-box adversarial attacks against Madry adversarial training [47] on a single, relatively simple ResNet50 model [31].
 The purpose of these experiments is to check to what extent black-box attacks, typically successful against vanilla models, remain as successful when confronted to adversarially trained models.
 This preliminary experiment provides us with the first insight into the impact of defenses studied in white-box settings on the effectiveness of black-box attacks.
-3.1 Experimental Setup
-We evaluated the attack success rate of 11 black-box adversarial attacks against a standardly trained model (vanilla) and an adversarially trained model.
+III-A Experimental Setup
 We selected a ResNet50 vanilla target and a ResNet50 robust target that underwent adversarial training.
-We opted for the same model architecture to eliminate the variations that could arise from differing model architectures.
-We followed the Robustbench evaluation protocol (Croce et al., 2020) and reused the original implementation of each attack.
-All attacks were untargeted and bound to an $L_{\infty}$ maximum perturbation with a distance of $4/255$.
-We evaluated the ASR on 5000 examples from the ImageNet validation set, we used the same image identifiers (IDs) as in RobustBench.
-All the experiments are repeated with three randoms seeds.
-As outlined in 2.1, we selected nine transfer-based and two query-based adversarial attacks to provide a comprehensive evaluation of adversarial techniques, balancing foundational, and SOTA approaches.
+We use the same model architecture to eliminate the variations that could arise from differing model architectures.
+We followed the Robustbench evaluation protocol [15] and reused the original implementation of each attack: All attacks are untargeted and bound to a $L_{\infty}$ maximum perturbation with a distance of $4/255$, unless mentioned otherwise.
+We evaluated the Attack Success Rate (ASR) on 5000 examples from the ImageNet validation set, and we used the same image identifiers (IDs) as in RobustBench.
+We selected nine transfer-based adversarial attacks and four query-based adversarial attacks to provide a comprehensive evaluation of adversarial techniques, balancing foundational approaches and SoTA.
 For transfer attacks, our selection emphasizes the diversity of techniques by including methods targeting input transformations (TI, DI), gradient optimization (MI, VNI, VMI, ADMIX), universality (UAP), architectural nuances (GHOST), and geometry (LGV).
-For query-based attacks, we focused on methods that improve query efficiency and performance by leveraging surrogate models, and that rely on different optimization techniques.
-BASES uses a small dimensional search space by modifying the weights given to each surrogate based on the feedback received from querying the target, while TREMBA learns a low-dimensional embedding then performs an efficient search within this space.
-As discussed in 2.1, all query attacks in this study require surrogate models.
-The hyperparameters for all transfer and query attacks are detailed in the Appendix A and B, respectively.
-We employed surrogate models that are either individually larger than the target model or collectively larger as an ensemble, allowing us to leverage the full capacity of the attack.
-For all single-surrogate attacks, including the five FGSM-based baseline attacks, ADMIX, UAP, GHOST, and LGV, the surrogate is a WideResNet50-2 (Zagoruyko, 2016) with standard training.
-For TREMBA, the surrogates are an ensemble of four models with standard training: VGG16 (Simonyan & Zisserman, 2014), ResNet18, SqueezeNet (Iandola et al., 2016), and GoogLeNet (Szegedy et al., 2015) as provided in their repository.
-For BASES, the surrogates are an ensemble of 10 models with standard training: VGG16BN, ResNet18, SqueezeNet11, GoogLeNet, MnasNet10 (Tan et al., 2019), DenseNet161 (Huang et al., 2017), EfficientNetB0 (Tan & Le, 2019), RegNetY400MF (Radosavovic et al., 2020), ResNeXt10132x8d (Xie et al., 2017), ConvNeXtSmall (Liu et al., 2022) from .
-The models used in this study were obtained from the torchvision library (maintainers & contributors, 2016) and the RobustBench benchmark Croce et al.
-(2020).
+For query-based attacks, we included two families.
+First, pure query attacks, which rely solely on feedback from the target model such that their performance is not influence by the choice of the surrogate model , such as (SQUARE) and (SIGNFLIP).
+Second, hybrid query attacks, which improve query efficiency and performance by leveraging surrogate models with optimization strategies to reduce the number of queries.
+Specifically, (BASES) modifies the weights assigned to each surrogate based on feedback from the target, effectively searching a small-dimensional space, while (TREMBA) learns a low-dimensional embedding and conducts an efficient search within this space.
+The details and hyperparameters for all attacks are provided in the Appendix -C.
+For all single-surrogate attacks, including the five FGSM-based baseline attacks, ADMIX, UAP, GHOST, and LGV, the surrogate is a WideResNet50-2 [81] with standard training.
+For TREMBA, the surrogates are an ensemble of four models with standard training: VGG16 [60], ResNet18, SqueezeNet [35], and GoogLeNet [64] as provided in their repository.
+For BASES, the surrogates are an ensemble of 10 models with standard training: VGG16BN, ResNet18, SqueezeNet11, GoogLeNet, MnasNet10 [65], DenseNet161 [32], EfficientNetB0 [66], RegNetY400MF [58], ResNeXt10132x8d [80], ConvNeXtSmall [46] .
+The models used in this study were obtained from the torchvision library [50] and the RobustBench benchmark [15].
 We excluded the SGM attack from this section because, at the time of this study, its implementation supports only ResNet and DenseNet surrogates, while our setup uses WideResNet50-2.
-3.2 Results
-We evaluate the ASR of the black-box attacks against standard and adversarially trained Resnet-50 models.
-We illustrate the mean ASR using bars, and the standard deviation (STD) as error bars in Figure 1.
-Against the undefended model, all attacks except UAP and LGV achieve more than 40% ASR, with the most recent attacks BASES and TREMBA reaching 81.08% ± 0.49 and 89.56% ± 0.09, respectively.
-However, all black-box attacks fail against the robustified model with an ASR of less than 4%.
-The most recent attacks only bring about marginal improvement over the simplest FGSM attacks.
-The most effective attack in our evaluation, TREMBA achieves 3.56% ASR while TI-FGSM reaches 2.81% ASR.
-To ensure that the reduced effectiveness of black-box attacks against the adversarially trained ResNet50 model is not due to budget limitation, we quadruple the budget of the best black-box attacks (BASES and TREMBA).
-Even with this increased budget, both attacks still significantly underperform against the adversarially trained ResNet50 compared to their performance on the vanilla ResNet50 model, as detailed in Appendix C.
-Figure 1: The blue bars (with error bars) show the success rates and standard deviations for the vanilla ResNet50 model, while the orange bars (with error bars) show the results for the robust ResNet50 model.
-Insight 1 The effectiveness of black-box attacks drastically decreases against a simple adversarially trained model, with the strongest attack seeing its ASR drops from 89.56% to 3.56%.
-4 Effectiveness of White-Box Defenses Against Black-Box Attacks
-Given that even the most recent advances in black-box attacks remain ineffective against Madry adversarial training, we investigate whether recent innovations in defense mechanisms (studied in white-box settings) further reduce the effectiveness of black-box attacks.
-In particular, we want to observe whether these defenses designed against white-box attacks (viz.
-AutoAttack) and established benchmarks (viz.
-Robustbench), generalize to black-box attacks.
-A positive answer to these questions would maintain (and even raise) the motivation to develop defenses in white-box settings, whereas a negative answer would raise an urgent necessity of investigating specific defenses against black-box attacks.
-4.1 Experimental Setup
-We followed the same protocol as in the previous section, in terms of datasets, and adversarial black-box attacks.
-We selected nine robust target models from the Robustbench leaderboard.
-We ensured to have an even distribution across the robustness spectrum for the nine targets.
-The robust accuracy of the most robust model, a Swin-L (Liu et al., 2024) is 59.64% (that is, a success rate of AutoAttack of 25.97%), and the robust accuracy of the least robust model, a ResNet18 (Salman et al., 2020) is 25.32% (for a success rate of AutoAttack of 50.53%).
-Table 1 presents the architecture and the number of parameters of the models.
-These nine models have been defended using different categories of mechanisms, listed in Table 2.
-We kept the same surrogate ensemble for BASES and TREMBA, and we used the surrogate model ResNet50 with standard training for single-surrogate attacks.
-In addition to the selected black-box attacks, we incorporated the SGM attack due to its compatibility with the ResNet50 surrogate model.
-Table 1: Summary of different models from the RobustBench leaderboard (Croce et al., 2020) and their configurations. Legend of the defenses in Table 2.
+III-B Results
+We evaluated the ASR of the black-box attacks against standard and adversarially trained Resnet-50 models.
+We illustrate the ASR using bars in Figure 1.
+Under a restrictive perturbation budget of 4/255, all attacks except UAP and SIGNFLIP achieve more than 40% ASR against the undefended model, with the most recent attacks LGV, BASES and TREMBA reaching 90.11% ± 0.12, 81.08% ± 0.49, and 89.56% ± 0.09, respectively.
+However, all black-box attacks fail against the robustified model with an ASR of less than 17%.
+To rule out query-budget limitations, we quadruple the budgets of the BASES and TREMBA hybrid attacks.
+Even with this increase, both attacks still perform far worse against the adversarially trained ResNet50 than against the vanilla model (see Appendix A-A).
+When the perturbation budget is increased to 16/255 and the query budget for query-only attacks is doubled, black-box attacks can achieve higher ASR.
+Under these settings, the performance gap between the undefended and defended models narrows, with the Square Attack reaching an ASR of 79%.
+Figure 1: The blue bars show the success rates and standard deviations for the vanilla ResNet50 model, while the orange bars show the results for the robust ResNet50 model.
+Bars in solid color are with attack distance $\epsilon=4/255$, while dashed bars are with attack distance $\epsilon=16/255$
+III-C Parameter Sensitivity Analysis
+We study in the following the impact of increasing perturbation and computation budgets.
+In Fig.2 we present the results for the increasing $\epsilon$ budgets, and report the exact performance with increased number of iterations and queries in Appendix B.
+Our results confirm that significantly increasing the perturbation budget does not lead to a collapse of robustness of adversarially trained models and that TREMBA is the only attack with a success rate increasing over 30% for extremely large budgets.
+The results in Table II show that increasing iterations has negligible impact, and Table III that increasing the number of queries has also negligible impact.
+Figure 2: Sensitivity study of the impact of the perturbation budget $\epsilon$.
+TABLE II: Parameter study of the impact of the iteration budget.
+Attack Iteration=N Iteration=2N Iteration=5N
+MI-FGSM 88.79 $\rightarrow$ 10.52 91.63 $\rightarrow$ 11.61 93.79 $\rightarrow$ 12.67
+DI-FGSM 94.55 $\rightarrow$ 12.11 97.32 $\rightarrow$ 13.30 99.17 $\rightarrow$ 14.38
+TI-FGSM 83.53 $\rightarrow$ 15.61 89.31 $\rightarrow$ 17.61 93.45 $\rightarrow$ 18.69
+VMI-FGSM 94.79 $\rightarrow$ 12.74 97.73 $\rightarrow$ 14.30 99.06 $\rightarrow$ 14.82
+VNI-FGSM 96.94 $\rightarrow$ 13.21 98.62 $\rightarrow$ 14.67 99.45 $\rightarrow$ 14.98
+ADMIX 98.28 $\rightarrow$ 14.24 98.90 $\rightarrow$ 14.58 99.43 $\rightarrow$ 15.02
+UAP 70.87 $\rightarrow$ 9.93 70.78 $\rightarrow$ 9.93 72.38 $\rightarrow$ 10.02
+GHOST 94.40 $\rightarrow$ 11.27 96.48 $\rightarrow$ 12.21 98.41 $\rightarrow$ 13.39
+LGV 99.66 $\rightarrow$ 15.16 99.92 $\rightarrow$ 16.61 99.97 $\rightarrow$ 16.39
+TABLE III: Parameter study of the impact of number of query budgets.
+Model Attack Query=N Query=2N
+Vanilla BASES 99.56 99.66
+Vanilla TREMBA 99.84 99.90
+Robust BASES 13.76 14.05
+Robust TREMBA 32.67 37.79
+Insight 1 The effectiveness of black-box attacks drastically decreases against a simple adversarially trained model, with the strongest attack (Square) seeing its ASR drop from 79.04% to 16.06%.
+IV Effectiveness of White-Box Defenses Against Black-Box Attacks
+Our first insight demonstrates that simple Madry adversarial training is an effective defense against most SoTA black-box attacks under a restrictive attack budget.
+We explore in this section whether improved defenses manage to also improve against SoTA black-box attacks in all the settings.
+These new defenses have been designed to overcome AutoAttack, the SoTA white-box attack, and we are the first to confront them to SoTA black-box attacks.
+IV-A Experimental Setup
+We followed the evaluation protocol of the first study on nine robust models from Robustbench.
+We selected models that were (1) peer-reviewed, (2) compatible with all the attacks (for instance Skip connections to run GHOST attacks), and (3) available in a diverse set of architectures (transformers , convolutions), size (Base and Large variants), and defense mechanisms (ensembling, synthetic data augmentation, downsampling layers).
+Our selection strategy allows to study the impact of defense mechanisms as a whole, and the impact of size and architecture within the same defense mechanism.
+We provide more details in Appendix -B.
+The most robust model, a Swin-L [43] achieves 59.64% robust accuracy, and the least robust model, a ResNet18 [59] achieves 25.32% robust accuracy.
+Table IV presents the architecture and the number of parameters of the models.
+These nine models have been defended using different mechanisms.
+Mechanism A involves adversarial training with extensive data augmentation, regularization, weight averaging, and extensive pretraining.
+Mechanism B involves non linear ensemble of two models, one defended with adversarial training and one vanilla model with standard training.
+Mechanism C mitigates adversarial examples using downsample convolutional layers before subsequent network layers.
+All models with standard adversarial training are denoted with defense D.
+TABLE IV: Robust models included in our evaluation and their configurations.
+Defense A: Adv training with large data augmentation, regularization, weights averaging, pretraining; Defense B: Non linear ensemble of two base models robust and vanilla; Defense C: Downsample convolutional layers before subsequent network layers; Defense D: Standard adversarial training
 Rank Reference Architecture Robust Accuracy Defense Architecture Type Parameters
-1 Liu et al. (2024) Swin-L (Liu et al., 2021) 59.56 % A Transformer 196.53M
-2 Bai et al. (2024) ConvNeXtV2+Swin-L (Liu et al., 2022; 2021) 58.50 % A, B Transformer & Convolution 394.49M
-3 Liu et al. (2024) ConvNeXt-L (Liu et al., 2022) 58.48 % A Convolution 197.77M
-5 Liu et al. (2024) Swin-B (Liu et al., 2021) 56.16 % A Transformer 87.77M
-7 Liu et al. (2024) ConvNeXt-B (Liu et al., 2022) 55.82 % A Convolution 88.59M
-12 Singh et al. (2024) ViT-S+ConvStem (Dosovitskiy et al., 2020; Xiao et al., 2021) 48.08 % C Transformer 22.78M
-17 Salman et al. (2020) WideResNet50.2 (Zagoruyko, 2016) 38.14 % D Convolution 68.88M
-18 Salman et al. (2020) Resnet50 (He et al., 2016) 34.96 % D Convolution 25.56M
-21 Salman et al. (2020) Resnet18 (He et al., 2016) 25.32 % D Convolution 11.69M
-Table 2: List of the defense mechanisms.
-Label Description
-A Adv training with large data augmentation, regularisation, weights averaging, pretraining
-B Non linear ensemble of two base models robust and vanilla
-C Downsample convolutional layers before subsequent network layers
-D Standard adversarial training
-4.2 Results
+1 [43] Swin-L [45] 59.56 % A Transformer 196.53M
+2 [5] ConvNeXtV2+Swin-L [46, 45] 58.50 % B Transformer & Convolution 394.49M
+3 [43] ConvNeXt-L [46] 58.48 % A Convolution 197.77M
+5 [43] Swin-B [45] 56.16 % A Transformer 87.77M
+7 [43] ConvNeXt-B [46] 55.82 % A Convolution 88.59M
+12 [61] ViT-S+ConvStem [24, 78] 48.08 % C Transformer 22.78M
+17 [59] WideResNet50.2 [81] 38.14 % D Convolution 68.88M
+18 [59] Resnet50 [31] 34.96 % D Convolution 25.56M
+21 [59] Resnet18 [31] 25.32 % D Convolution 11.69M
+IV-B Results
 Figure 3: Success rate of blackbox attacks against SoTA defenses.
-We analyze the performance of defenses against black-box attacks by examining the impact of different defense strategies, size of models, and architecture of models, as well as the relationship between AutoAttack success rates and black-box attacks success rates.
-Impact of defense mechanism.
-We categorize the defended models from Table 1 into four distinct families (A, B, C, and D), reflecting the various training strategies employed to enhance model robustness.
-In Figure 2 we show the ASR of the 12 black-box attacks against the four defenses (robust models).
-The models are sorted according to the Robustbench leaderboard, with stronger defenses against AutoAttack positioned to the right.
-Models using the same category of defense mechanisms but differing architectures are grouped in the same color.
-We observe that stronger defenses against AutoAttack lead to stronger defenses against black-box attacks.
-For instance, Defense A (red) outperforms Defense C (yellow), which in turn surpasses Defense D (blue) across all black-box attacks.
-An exception to this trend is observed with the ensemble Conv+Swin Defense, ranked #2 on Robustbench and mixing mechanisms of types A and B.
-Although ranked second against AutoAttack, it performs similarly to Defense D when subjected to black-box attacks.
-This is explained by the composition of defense mechanisms A and B, which combines the adversarially trained model from Defense A with a standard (vanilla) model, as explained in Section 2.2.
-This hybrid nature introduces a vulnerability: while this defense ranks among the best defenses against AutoAttack, the inclusion of a standard model component causes its performance against black-box attacks to align with that of weaker AutoAttack defenses.
-This vulnerability arises because black-box attacks often exploit vanilla surrogates, making the defense susceptible due to the standard model’s presence.
-Table 3: Success rate on Resnet models trained by Salman et al. (2020).
-Attacks Salman ResNet18 Salman ResNet50 Salman WideResNet50.2
-AUTOATTACK 50.53 43.84 42.87
+We study the performance of defenses against black-box by considering the impact of different defense strategies, the size of the models, and the architecture of the models.
+We also analyze these results by discussing the relationship between AutoAttack success rates and black-box attack success rates.
+Impact of defense mechanism
+We report in Figure 3 the ASR of the 14 black-box attacks against each of the defended models.
+The models are sorted from left to right according to the Robustbench leaderboard, with stronger defenses against AutoAttack on the right.
+Models using the same category of defense mechanisms are grouped with the same color.
+Defenses from the family A (red) outperform all other defenses.
+Defense C (orange) is better than ensembling (green), and vanilla adversarial training (blue).
+Surprisingly, the Conv + Swin ensemble model (orange, ranked 2 against AutoAttack) significanly underperforms models supposedly ”weaker” according to Robustbench, and is often weaker than old defenses using only standard training.
+Conv + Swin is weaker than the vanilla adversarial models against MI-FGSM, DI-FGSM, VM-FGSM, VNI-FGSM, SGM, GHOST, and BASES.
+We hypothesize that this vulnerability arises because black-box attacks often exploit vanilla surrogates, making the ensemble defense susceptible due to the presence of the standard model.
+These trends remain consistent when increasing the epsilon values from $4/255$ to $8/255$ and $16/255$, as well as when extending the iteration budget from N to 2N and 5N.
+Similarly, increasing the query budget from $N$ to $2N$ does not alter the observed pattern.
+The detailed results are presented in Appendix B.
+TABLE V: Success rate on Resnet models trained by Salman et al.[59].
+Attacks ResNet18 ResNet50 WideResNet50.2
+AA 50.53 43.84 42.87
 MI-FGSM 1.28±0.00 1.68±0.00 1.34±0.00
@@
 GHOST 1.21±0.18 1.69±0.05 1.49±0.06
-LGV 1.63 ± 0.13 1.72 ± 0.26 1.52 ± 0.08
+LGV 2.27±0.11 2.92±0.01 2.58±0.09
 BASES 2.36±0.03 2.83±0.04 2.35±0.06
 TREMBA 4.06±0.06 3.56±0.06 3.22±0.06
-Impact of model size.
-We compare in Table 3 the success rates of the black-box attacks against similar adversarially trained models.
+SQUARE 20.58±0.21 16.05±0.21 15.09±0.1
+SIGNFLIP 4.6±0.06 4.29±0.1 3.79±0.06
+Impact of model size and architecture
+We focus now on models using the same defense mechanism, and compare the impact of the size of model and archtitectures.
+We compare in Table V the success rate of the black-box attacks against models defended with standard adversarial training (family D).
 These models do not benefit from advanced defense mechanisms (e.g., synthetic data augmentation, …) and only differ in size.
-While larger models with adversarial training tend to achieve higher robustness to AutoAttack, the increase in size does not lead to lower success rates.
-Similar insights are revealed in Table 4, where we compare the robustness of two sizes of ConvNext models and SWIN Transformers.
-Larger models do not consistently demonstrate stronger robustness to adversarial perturbations in black-box settings, although increased size improves robustness against AutoAttack.
-Impact of model architecture.
-Liu et al.
-(2024) demonstrate that the best robustness against AutoAttack can be achieved using indiscriminately Convnets or Transformers models.
-Our results in Table 4 confirm that for a fixed size, both architectures demonstrated similar robustness against black-box attacks.
-Table 4: Success rate on very large models trained by Liu et al. (2024).
-Attacks Liu ConvNeXt-B Liu Swin-B Liu ConvNeXt-L Liu Swin-L
-AUTOATTACK 28.01 27.14 25.95 25.98
+While scaling up the models from ResNet18 to WideResNet50 lowers AutoAttack (AA) success from $50.53\%$ to $42.87\%$, this pattern does not hold always for black-box attacks.
+The black-box attacks that suffer from the increase in target size are TREMBA, SQUARE, and SIGNFLIP with the SQUARE attack dropping from $20.58\%$ to $15.09\%$.
+Moving to Defense A trained models (extensive data enhancement, much larger models, and longer trainings), we report the impact of the increase in model size in Table VI.
+While increasing the size of the model (ConvNeXt-B $\rightarrow$ ConvNeXt-L, Swin-B $\rightarrow$ Swin-L) decreases the ASR of AutoAttack, it has no effect on black-box attacks.
+SIGNFLIP ASR marginally decreases from $1.63\%$ to $1.42\%$ on convolution models with increased size, but marginally improves from $1.28\%$ to $1.65\%$ on Swin transformers with increased size.
+TABLE VI: Success rate on very large models trained by Liu et al.[43].
+Attacks ConvNeXt-B Swin-B ConvNeXt-L Swin-L
+AA 28.01 27.14 25.95 25.98
 MI-FGSM 0.63±0.00 0.37±0.00 0.49±0.00 0.56±0.00
@@
 GHOST 0.63±0.03 0.46±0.07 0.49±0.09 0.49±0.04
-LGV 0.57 ± 0.04 0.36 ± 0.05 0.45 ± 0.05 0.46 ± 0.04
+LGV 0.92±0.07 0.91±0.03 1.07±0.03 1.03±0.03
 BASES 1.38±0.06 0.96±0.02 1.16±0.00 1.26±0.02
 TREMBA 1.34±0.04 1.14±0.03 1.47±0.01 1.25±0.01
-Relationship between AutoAttack success rate and black-box success rate.
-As shown in Figure 3, the scatter plot illustrates the relationship between the ASR of AutoAttack and that of 12 black-box attacks.
+SQUARE 8.9±0.1 8.71±0.12 8.56±0.04 8.65±0.11
+SIGNFLIP 1.63±0.06 1.28±0.03 1.42±0.03 1.65±0.12
+Figure 4: Relation between success rate of AutoAttack and the success rate of black-box attacks.
+Relationship between AutoAttack success rate and black-box success rate
+As shown in Figure 4, the scatter plot illustrates the relationship between the ASR of AutoAttack and that of 14 black-box attacks.
 Each data point represents the ASR on the y-axis of a given black-box attack (indicated by a color) against a specific robust model (determined by its AutoAttack success rate on the x-axis).
-Although the initial ASR of black-box attacks is low (less than 5%), we observe a clear downward trend as robust models improve their effectiveness against AutoAttack, until the best four robust models (with AutoAttack success rate ranging from 25.48% to 27.14%).
-This vulnerability stems from the lack of correlation between model size and robustness, as these models all employ defense (A) with diverse sizes and architectures, as shown in Table 1.
+We observe a clear downward trend as robust models improve their effectiveness against AutoAttack, until the best four robust models (with AutoAttack success rate ranging from 25.48% to 27.14%).
+This vulnerability stems from the lack of correlation between model architecture and robustness, as these models all employ defense (A) with diverse sizes and architectures, as shown in Table IV.
 Furthermore, the best robust model (25.48% AutoAttack success rate) underperforms compared to the other top robust models against black-box attacks.
 This occurs because it uses the ensemble defense (B), which, as discussed earlier, combines a robust model with a standard model, making it vulnerable to black-box attacks that often exploit surrogates resembling standard models.
-Figure 3: Relation between success rate of AutoAttack and the success rate of black-box attacks.
-Insight 2 SoTA defenses against AutoAttack generally correlate with improved black-box robustness.
-However, having a large model size does not always guarantee better robustness.
-Additionally, an ensemble defense that achieves similar robustness to a single defense against AutoAttack may be less effective against black-box attacks, especially those using surrogate models similar to the ensemble’s components.
-5 Robust surrogates to Improve Black-Box Attacks
+Insight 2 SoTA defenses improve robustness to black-box attacks.
+However, larger models do not guarantee higher robustness and effective ensemble defense against AutoAttack may be less effective against black-box attacks, especially those using surrogate models similar to the ensemble’s components.
+Large data augmentation is the best defense against black-box attacks.
+V Robust surrogates to Improve Black-Box Attacks
 The results from the previous section demonstrated that models robust against AutoAttack remain relevant (to some extent) against black-box attacks.
-A critical concern then is whether these robust "defenses" could not be used by the attacker to improve its black-box attack.
+A critical concern then is whether these robust ”defenses” could not be used by the attacker to improve its black-box attack.
 Indeed, most of the black-box require designing a surrogate and can therefore leverage the recent advances in model defense to strengthen their attacks.
 We hypothesize that these new defenses raise security risks by providing stronger surrogates to attackers.
-5.1 Experimental Setup
-We adhered to the same protocol as in previous sections, maintaining consistency in datasets and adversarial black-box attacks.
-We considered various configurations regarding the robustness of both the surrogate and target models.
-For the target models, we selected highly robust models (Swin-L, ConvNeXtV2+Swin-L, ConvNext-L from (Liu et al., 2024) and (Bai et al., 2024)), a moderately robust model (ResNet18 from (Salman et al., 2020)), and a non-robust model (ResNet50).
+V-A Experimental Setup
+We adhered to the same protocol as in previous sections.
+We considered various configurations of surrogate and target models.
+For the target models, we selected highly robust models (Swin-L, ConvNeXtV2+Swin-L, ConvNext-L from [43] and [5]), a moderately robust model (ResNet18 from [59]), and a non-robust model (ResNet50).
 This way, we can measure the effects of using robust surrogates on targets with top defense (A), a simple defense (D) and no defense.
-As for surrogate, we used robust and non-robust models.
-The selection of robust surrogates was based on choosing the highly ranked robust models that are not used as target and are compatible with the attack methods.
-Hence, we used the robust surrogate RaWideResNet-101-2 from (Peng et al., 2023) for single surrogate attacks, because it is the best defense in Robustbench that supports skip connections (which are essential for the GHOST attack).
-For ensemble surrogate attacks, we also added the robust surrogate Swin-B from (Liu et al., 2024), as it is the best robust model available, excluding the target models.
+We used the robust surrogate WideResNet-101-2 from [54] for single surrogate attacks, because it is the best defense in Robustbench that supports skip connections (which are essential for the GHOST attack).
+For ensemble surrogate attacks, we also added the robust surrogate Swin-B from [43], as it is the best robust model available, excluding target models.
 For single, non-robust surrogate attacks, we used a vanilla WideResNet-101-2 in order to have the same architecture as in the robust surrogate, ensuring fair comparison.
-For BASES and TREMBA (who use an ensemble of surrogates), we kept the same ensemble of vanilla models as in our previous experiments.
+For BASES and TREMBA (which use an ensemble of surrogates), we kept the same ensemble of vanilla models as in our previous experiments.
 We excluded the SGM attack that does not support the WideResNet-101-2 surrogates used in our setup.
-5.2 Results
-We present in Figure 4 the success rate of black-box attacks using non-robust surrogates and robust surrogates against a non robust target and robust targets.
+V-B Results
+We present in Figure 5 the success rate of black-box attacks using non-robust surrogates and robust surrogates against a non-robust target and robust targets.
 Figure 5: Success rates of black-box attacks using vanilla and robust surrogates against a vanilla target and robust targets.
-Non-robust target: The results clearly show that using robust surrogate models is detrimental to attack success rates.
-Across all black-box attacks, non-robust surrogates consistently outperform robust surrogates, with an average improvement of 35.39 percentage points in attack success rates on the vanilla ResNet50 model.
+Non-robust target: The results clearly show that the use of robust surrogate models is detrimental to the success rates of the attack.
+Across all black-box attacks, non-robust surrogates consistently outperform robust surrogates, with an average improvement of 35.19 percentage points in attack success rates on the vanilla ResNet50 model.
 For instance, GHOST improves from 17.95% to 49.36%, BASES from 17.94% to 81.08%, and TREMBA from 67.54% to 89.56%.
 Even UAP that had the lowest success rate improves from 0.73% to 7.11%.
-Robust target: By contrast, robust surrogates yield significant improvement in the success rate of black-box attacks when targeting robust models.
+Notably, attacks that leveraged a training phase before the attacking phase (LGV and TREMBA) exhibited a relatively smaller improvement gap compared to other attacks as robust surrogates were already performing well.
+Specifically, LGV improved from 81.34% to 87.39%, while TREMBA increased from 67.54% to 89.56%.
+Robust target: By contrast, robust surrogates yield significant improvement in the success rate of black-box attacks when targeting robust models, resulting in an average improvement of 6.42 percentage points in ASR across attacks and target models.
 Notably, attacks using robust surrogates on top-tier robust models (e.g., Liu ConvNeXt-L and Liu Swin-L) yield up to 15 times higher success rates than their non-robust counterparts.
@@
 Similarly, BASES improves its success rate from 1.26% to 12.55% against the Swin-L model.
-The significant improvement of success rate is consistent across all defense mechanisms, and all attacks, except LGV, and UAP.
-We hypothesize that the initial training phase of the LGV models, where a high learning rate is used over additional epochs, potentially weakens the model’s robustness during the early sampling phase, which contrasts with GHOST networks that maintain stronger robustness after their creation.
-This hypothesis is explored in Appendix D.
+An improvement of more than or equal to 2 percentage points in ASR is consistent across all defense mechanisms and all attacks, except LGV and UAP.
+We hypothesize that the initial training phase of the LGV models, where a high learning rate is used over additional epochs, weakens the model’s robustness during the early sampling phase, which contrasts with GHOST networks that maintain stronger robustness after their creation.
+We explore this hypothesis in Appendix B-A.
 The results highlight that the choice of surrogate model plays a critical role in the effectiveness of transfer-based black-box attacks, particularly in restricted environments where the robustness of the target is unknown.
@@
 As a result, transfer attacks on robust surrogates aim to fool other regions of the input space that are not yet secured by adversarial training, making them better suited for attacking robust models.
-On the other hand, transfer attacks on vanilla surrogates focuses more on blind spot regions of the input space that are also likely to be mispredicted by other vanilla target models.
-Insight 3 The choice of the surrogate model, robust or vanilla, impacts the success of black-box attacks.
+On the other hand, transfer attacks on vanilla surrogates focus more on blind spots of the input space that are also likely to be mispredicted by other vanilla target models.
+Insight 3 The choice of the surrogate model, robust or vanilla, affects the success of black-box attacks.
 Robust surrogates are more effective against robust targets, while vanilla surrogates excel against vanilla targets.
 Attackers can therefore improve attack effectiveness by adaptively selecting surrogates based on the perceived robustness of the target model.
-6 Conclusion and Perspectives
+VI Discussion
+VI-A Impact of defense mechanisms on different attack strategies
+Against advanced defenses, all attack mechanisms are disrupted both with $\epsilon=4/255$ and with $\epsilon=16/255$ (9), meanwhile some attacks mechanism remain effective under standard adversarial training (SAT): Compared to vanilla models, SAT is sufficient against all our baseline attacks under low $\epsilon=4/255$ budgets.
+Under higher $\epsilon=16/255$, only attacks that leverage the decision boundary similarity mechanism (Signflip/Square) and the knowledge transfer hypothesis (TREMBA) achieve over 20% success rate.
+We explain this observation by the lower impact of SAT on these two mechanisms compared to the other mechanisms.
+Indeed, SAT alone is not sufficient to disrupt the decision boundaries enough when comparing two models with the same architectures, and the same training data (same ”knowledge”).
+The data augmentations in defenses A and the ensembles in defenses B, however, disrupt TREMBA, Square and Signflip by shifting the knowledge and the boundaries.
+Attacks like SQUARE, which rely on querying the boundary to estimate its position, struggle because the boundary geometry of defenses A becomes fundamentally different from the surrogate’s.
+Meanwhile, the gradient similarity attacks like MI-FGSM and VMI-FGSM rely on the assumption that the gradient direction is stable and transferable, which is disrupted by SAT (by smoothing the loss landscape).
+Using extensive data augmentation and model scaling (defenses A), the boundaries, the knowledge and gradients are ”decorrelated” compared to the source model.
+These defenses also disrupt shared space attacks (DI, GHOST, LGV, UAP) that exploit large, contiguous volumes of empty space in the data manifold to generate adversarial - the large-scale data augmentation fill these spaces, and prevent UAP from finding universal directions of attacks.
+Interaction-based attacks (TI-FGSM) and non-robust feature attacks (ADMIX) try to find shortcuts, using respectively low-order (pixel-local) interactions and non-robust features.
+Our results confirm that these shortcuts are fundamentally disrupted but all adversarial training defenses, making them unreliable for transfer attacks.
+VI-B Impact of Adaptive Attacks.
+Our third insight confirms that implementing adaptive attacks by including a robust surrogate when crafting the adversarial examples may not be a suitable approach.
+Although it improves the effectiveness of the attacks (up to 6.42%), the final success rate achieves at best 12.5% using adaptive BASES against the strongest defense (Liu Swin-L).
+Meanwhile, adapting the attacks using a surrogate with a specific defense decreases the effectiveness of the attack against vanilla models.
+Therefore the adaptation needed in the black-box setting should be carefully crafted for each target.
+To better understand this phenomenon, we measure the gradient similarity across 1,000 ImageNet samples.
+For each image-label pair, we calculate the gradient of the cross-entropy loss with respect to the input image for both surrogate and target models.
+We report the cosine similarity between the gradients.
+This process was repeated for all surrogate-target model pairs, and the similarities were averaged across all samples to produce a similarity matrix in Fig.
+6.
+Each cell in the matrix represents the average cosine similarity between a surrogate model (rows) and a target model (columns), with values ranging from 0 (orthogonal gradients) to 1 (perfectly aligned gradients).
+Our results demonstrate that robust targets have a higher gradient alignment with robust surrogates than with vanilla surrogates, even when robust surrogates are from very different architectures and sizes.
+Figure 6: Gradient similarity between surrogates and targets over 1000 samples.
+VI-C Transferability Bounds Under Mild Robustness Mechanisms
+Our study focused on SoTA defenses, that have all been designed against strong attacks ($\epsilon=8/255$ in Robustbench).
+We did not explore the impact of transferability when the models are midly robustified.
+However, the work promoting mild robustification [82, 83] was justified only by the tradeoff between robustness and accuracy that was ovserved then.
+Today, however, very large models and extensive data augmentation allows to robustify the models with agressive attacks and strong mechanisms without degradation of their clean performance (quite the opposite [73]).
+VII Limitations and Perspectives
+While our work investigated a large set of black-box attacks and robust models, we mostly focused on extremely robust models.
+[84] for instance focused on mildly perturbed examples and uncovered their interactions with simple adversarial training.
+Expanding our study of transfer-based attacks to consider varying levels of attack strength and robustness levels between surrogate and target models would also help to deepen our understanding of transferability dynamics.
+In particular, considering randomness in the pre-processing, and defenses including obfuscation and randomness.
+In addition, while this study focused on image classification tasks, future research could extend the analysis to other tasks, such as object detection or segmentation, to understand how black-box robustness may vary across applications.
+These regression tasks raise a new challenge of defining acceptable robustness thresholds, given that the output is not binary.
+Although there is an established benchmark for adversarial robustness for classification tasks, there is no such benchmark for segmentation tasks.
+Another limitation is that while we included large transformers models, we did not explore the robustness of foundations models such as GPT [57], LLaMA [67], CLIP [56], and SAM [39], or their variants.
+Research on robustification of these models is still in its infancy for a benchmark.
+VIII Conclusion
 This research investigated the effectiveness of black-box adversarial attacks against SoTA and standard robust models using ImageNet, structured around three series of experiments, uncovering key insights..
@@
 Consequently, understanding the limitations of current black-box attacks against robust models is crucial for developing more effective attack strategies.
-Based on our second experiments, we found that defense mechanisms optimized against AutoAttack often generalize effectively to resist black-box attacks, even though performance differences are less pronounced in black-box settings.
+Based on our second experiment, we found that defense mechanisms optimized against AutoAttack often generalize effectively to resist black-box attacks, even though performance differences are less pronounced in black-box settings.
 This finding supports the relevance of AutoAttack as a benchmark for black-box robustness and suggests that current efforts focused on white-box defenses provide a meaningful degree of robustness in black-box scenarios as well – though ensemble defenses may not always translate their AutoAttack robustness to black-box robustness, especially against attacks using similar surrogate models as the ensemble.
 Finally, we observed that selecting a surrogate with a level of robustness aligned with that of the target can significantly impact the success rate of transfer-based attacks.
-This finding reveal the possibility for strategic selection of surrogate models when planning black-box attacks.
-Future work should consider designing black-box attacks that specifically target SOTA robust models to better reflect real-world defensive capabilities.
+This finding reveals the possibility for strategic selection of surrogate models when planning black-box attacks.
+Future attacks should consider designing black-box attacks that specifically target SOTA robust models to better reflect real-world defensive capabilities.
 Additionally, further exploration of why SoTA defenses optimized for AutoAttack are also relevant for black-box robustness could yield valuable insights into cross-attack resilience mechanisms.
-Expanding the study of transfer-based attacks to consider varying levels of robustness between surrogate and target models would also help to deepen our understanding of transferability dynamics.
-Finally, while this study focused on image classification tasks, future research could extend the analysis to other tasks, such as object detection or segmentation, to understand how black-box robustness may vary across applications.
-In conclusion, this study demonstrates that SoTA defenses are notably resilient against black-box attacks, underscoring the importance of developing more targeted attack strategies to effectively challenge these modern robust models.
+In conclusion, our study demonstrates that SoTA defenses are notably resilient against black-box attacks, underscoring the importance of developing more targeted attack strategies to effectively challenge these modern robust models.
 As the field advances, these insights can guide both the creation of more sophisticated black-box attacks and the design of improved defense mechanisms that support the deployment of robust AI systems across a wide range of applications.
-Acknowledgment
-This research is supported by the Luxembourg National Research Fund under the AFR grant. We extend our gratitude to Martin Gubri (ParameterLabs) for his invaluable support with the LGV implementation.
-Appendix
-Appendix A Hyperparameters of Transfer Attacks
-Tables 5 and 6 provide a comprehensive overview of the hyperparameters used for the ten transfer-based adversarial attacks in our experiments.
-They categorize the key hyperparameters into two groups: common hyperparameters shared across most attacks and attack-specific hyperparameters that define the unique characteristics of each method.
-Most of the attack hyperparameters were kept at their default values, as specified in the repositories or libraries that provide the implementations of adversarial attacks (Kim, 2020; Woo, 2024).
-A.1 Common Hyperparameters
-Table 5 provides the default values and explanations for these common hyperparameters, such as the step size (Alpha), momentum factor (Decay), and number of iterations (Steps), which are essential for iterative attacks as they control the perturbation magnitude and the optimization process.
-Table 5: Common hyperparameters for transfer-based attacks.
-PARAMETER DEFAULT VALUE EXPLANATION APPLIES TO
-Alpha 2/255 Step size or update rate for perturbation. All attacks except UAP
-Decay 1.0 Momentum factor. All attacks except UAP
-Steps 10 Number of iterations. All attacks except UAP
-Epsilon (eps) 4/255 Maximum perturbation for adversarial example. All attacks
-A.2 Attack-Specific Hyperparameters
-In addition to the shared hyperparameters, each attack incorporates its own unique hyperparameters tailored to its specific mechanisms.
-Table 6 summarizes these attack-specific hyperparameters, listing the default values, explanations, and the corresponding attacks to which they apply.
-Table 6: Attack-specific hyperparameters for transfer-based attacks.
-PARAMETER DEFAULT VALUE EXPLANATION APPLIES TO
-Resize Rate 0.9 Rate at which the input images are resized. DIFGSM, TIFGSM
-Diversity Prob 0.5 Probability to apply input diversity. DIFGSM, TIFGSM
-Kernel Name Gaussian Type of kernel used. TIFGSM
-Kernel Length 15 Length of the kernel. TIFGSM
-Sigma (nsig) 3 Radius of the Gaussian kernel. TIFGSM
-N 5 Number of samples in the neighborhood. VMI, VNI
-Beta 1.5 Upper bound for neighborhood. VMI, VNI
-Prior Type no_data Type of prior used for attack optimization. UAP
-Patience Interval 5 Number of iterations to wait before checking convergence. UAP
-Gamma 0.5 Decay factor for gradients from skip connections. SGM
-LGV Epochs 5 Number of epochs for training the LGV models. LGV
-LGV Models Epoch 2 Number of models used per epoch. LGV
-LGV Learning Rate 0.05 Learning rate for the LGV training phase. LGV
-Randomized Modulating Scalar 0.22 Drawn from the uniform distribution 1-scalar, 1+scalar. GHOST
-Portion 0.2 Portion for the mixed image. Admix
-Size 3 Number of randomly sampled images. Admix
-Appendix B Hyperparameters of Query Attacks
-Tables 7 and 8 provides an overview of the hyperparameters used in BASES and TREMBA, respectively.
-Most of these hyperparameters were kept at their default values, as specified in their repositories, since our experiments aimed to evaluate their performance under standard configurations.
-For BASES, key hyperparameters include the ensemble size (N_wb), the type of fusion for surrogate models (Fuse), and the number of weight updates (Iterw).
-Similarly, TREMBA, which operates in two phases (training and attacking), relies on hyperparameters such as the number of training epochs (Epochs), the number of attack iterations (Num_Iters), the number of samples for Natural Evolution Strategy (NES), and latent space parameters like the Gaussian noise standard deviation (Sigma).
-Table 7: Hyperparameters for the BASES attack.
-PARAMETER DEFAULT VALUE EXPLANATION
-Epsilon (eps) 4 Perturbation bound magnitude (0 to 255 pixel range)).
-N_wb 10 Number of models in the ensemble.
-Bound linf Perturbation bound norm type.
-Iters 10 Number of inner iterations.
-Fuse loss Fuse method, e.g., loss or logit.
-Loss_name CW Loss function used for optimization.
-Algo PGD White-box algorithm used for the perturbation machine
-X 3 Factor to scale the step size alpha.
-Learning Rate (lr) 0.005 Learning rate for weight updates.
-Iterw 20 Number of iterations to update weights.
-Table 8: Hyperparameters for the TREMBA attack.
-PARAMETER DEFAULT VALUE EXPLANATION
-Epsilon (eps) 8/255 (Training Phase), 4/255 (Attacking Phase) Perturbation bound magnitude.
-Learning Rate (G) 0.01 (Training Phase), 5.0 (Attacking Phase) for gradient updates, for latent updates
-Momentum 0.9 (Training Phase), 0.0 (Attacking Phase) for SGD, for attack on the embedding space.
-Num_images 49k (Training Phase) Size of the training set
-Epochs 500 Number of epochs for Training.
-Schedule 10 Epochs between Training learning rate decay.
-Gamma 0.5 Training Learning rate decay factor.
-Margin 200.0 (Training Phase), 5.0 (Attacking Phase) Margin for the loss function.
-Sample_Size 20 Number of samples for NES
-Num_Iters 200 Number of iterations for NES.
-Sigma 1.0 Standard deviation of the Gaussian noise for NES.
-Learning Rate Min 0.1 Minimum learning rate in the Attacking Phase.
-Learning Rate Decay 2.0 Learning rate decay factor in the Attacking Phase.
-Plateau_Length 20 Number of iterations to monitor the objective function
-Plateau_Overhead 0.3 Allowed loss threshold to change before the decay
-Appendix C Increasing the Budget of Black-Box Attacks
-We quadrupled (multiplied by 4) the budget of the best black-box attacks in section 3 (BASES and TREMBA) against the robust ResNet50 model, to reassure that black-box attacks reduced effectiveness (compared to their effectiveness against the vanilla ResNet50 model) is not due to a limited budget.
+LLM usage considerations
+LLMs were used for editorial purposes in this manuscript, and all outputs were inspected by the authors to ensure accuracy and originality.
+Similarly, LLMs were used to assist in the process of writing code, and searching for relevant publications and literature reviews.
+All code and publications were manually verified and tested to ensure coherence in the results.
+Acknowledgements
+This research was funded in whole, or in part, by the Luxembourg National Research Fund (FNR), grant reference 18886361.
+Dr.
+Ghamizi is supported by the Luxembourg National Research Fund (FNR) CORE C24/IS/18942843.
+Appendix A Computational Budget of Attacks
+Table XIII details the computational budgets of all studied black-box attacks, quantified as the number of model calls required during the attacking phase per entry.
+TABLE XIII: Computational budgets of black-box attacks based on model calls
+Attack Algorithm Number of Model Calls Explanation
+MIFGSM Iterations (e.g., 10) MIFGSM iteratively refines the adversarial image over a number of steps,
+each requiring a model call
+DIFGSM Iterations (e.g., 10) DIFGSM performs iterative updates with input diversity applied at each step,
+requiring a model call per step
+TIFGSM Iterations (e.g., 10) TIFGSM performs iterative updates with translation-invariant gradient smoothing
+at each step, requiring a model call per step
+VMIFGSM Iterations * (N + 1) (e.g., 60) VMIFGSM iterates Steps times. In each iteration, there’s one model call for the adversarial
+image and N model calls for the neighboring images to calculate the gradient variance
+VNIFGSM Iterations * (N + 1) (e.g., 60) VNIFGSM, similar to VMIFGSM. In each iteration, there is one model call for the Nesterov and
+N model calls for the neighboring images to calculate the gradient variance
+GHOST Iterations (e.g., 10) GHOST applies Gaussian noise to the skip connection during each iteration
+(one Ghost model per iteration), requiring one model call per step
+LGV Iterations (e.g., 10) LGV uses an ensemble of models (lgv_nb_models_epoch * lgv_epochs models),
+with one model call at each step for each model in the ensemble
+Admix Iterations * Scale * Size (e.g., 150) In each step, Admix generates Size admixed images, scales them Scale times,
+and then calculates the loss
+UAP Max Iterations + Evaluation Calls The primary loop runs for max_iter iterations. Additionally, a function is
+(preparing the perturbation), then 1 called periodically to evaluate the current perturbation
+TREMBA (num_iters + 1) * sample_size (e.g., 4000) Tremba iteratively refines the adversarial perturbation within the latent space of a generator network.
+Each iteration involves generating sample_size perturbations and evaluating their losses
+BASES iterw * n_iters * n_wb (e.g., 2,000) BASES uses a surrogate ensemble containing n_wb models and performs an attack
+for n_iters iterations during each query (iterw)
+A-A Increasing the Computation Budget of Black-Box Attacks
+We are keeping the perturbation budget $\epsilon=4/255$, and quadrupled (multiplied by 4) the budget of the best black-box attacks in section III (BASES and TREMBA) against the robust ResNet50 model, to reassure that black-box attacks reduced effectiveness (compared to their effectiveness against the vanilla ResNet50 model) is not due to a limited budget.
 For BASES, we increased the number of inner iterations of PGD within the perturbation machine (Iters) to 40.
@@
 Similarly, TREMBA’s ASR is at 5.24% on the robust ResNet50 Budget (x4) compared to 89.56% on the vanilla ResNet50.
-These findings confirm that our earlier results in section 3 were not a consequence of budget limitations.
+These findings confirm that our earlier results in section III were not a consequence of budget limitations.
 Figure 7: The blue bars show the success rate for the vanilla ResNet50 model, while the orange bars show the results for the robust ResNet50 model, and the green bars show the success rate for the ResNet50 model by quadrupling the budget of the attacks.
-Appendix D AutoAttack Robustness of LGV and GHOST Models
+Appendix B Increasing the Perturbation Budget of Black-Box Attacks
+We have increased the perturbation budget to cover $\epsilon\in\{4/255,8/255,16/255\}$. In Table XIV, Table XV, and Table XVII, we only change the perturbation budget $\epsilon$.
+Once we have increased the perturbation budget to $\epsilon=16/255$, we first study the impact of increasing the iteration budget in Table II, and Table XVI, then, we study the impact of increasing the query budget in Table III, and Table XVIII.
+To further examine the impact of different defense strategies with increasing epsilon values and budgets, we present the results in the form of figures.
+Specifically, Figure 8 and Figure 9 correspond to epsilon values of 8/255 and 16/255, respectively.
+Additionally, Figure 10 for iteration and query budget of 2N and Figure 11 illustrate results for iteration budgets of 5N.
+TABLE XIV: Attack Performance at Different Epsilon Values.
+Attack Epsilon 4/255 Epsilon 8/255 Epsilon 16/255
+MI-FGSM 57.77 $\rightarrow$ 1.53 77.33 $\rightarrow$ 3.53 88.79 $\rightarrow$ 10.52
+DI-FGSM 71.31 $\rightarrow$ 1.94 87.07 $\rightarrow$ 4.18 94.55 $\rightarrow$ 12.11
+TI-FGSM 43.59 $\rightarrow$ 2.30 66.35 $\rightarrow$ 4.68 83.53 $\rightarrow$ 15.61
+VMI-FGSM 66.79 $\rightarrow$ 1.66 87.59 $\rightarrow$ 3.97 94.79 $\rightarrow$ 12.74
+VNI-FGSM 68.00 $\rightarrow$ 1.85 90.10 $\rightarrow$ 4.09 96.94 $\rightarrow$ 13.21
+ADMIX 77.90 $\rightarrow$ 2.54 93.04 $\rightarrow$ 4.68 98.28 $\rightarrow$ 14.24
+UAP 8.89 $\rightarrow$ 0.98 25.70 $\rightarrow$ 2.37 70.78 $\rightarrow$ 9.93
+GHOST 64.53 $\rightarrow$ 1.68 83.87 $\rightarrow$ 3.59 94.40 $\rightarrow$ 11.27
+LGV 90.11 $\rightarrow$ 1.33 98.07 $\rightarrow$ 5.25 99.66 $\rightarrow$ 15.16
+BASES 81.08 $\rightarrow$ 2.83 96.06 $\rightarrow$ 5.09 99.56 $\rightarrow$ 13.76
+TREMBA 89.56 $\rightarrow$ 3.56 99.45 $\rightarrow$ 9.52 99.84 $\rightarrow$ 32.67
+TABLE XV: Transfer Attack Performance Across Increased Epsilon 4/255, 8/255 and 16/255.
+Model MI-FGSM DI-FGSM TI-FGSM VMI-FGSM VNI-FGSM ADMIX UAP SGM GHOST LGV
+ResNet18 1.28 $\rightarrow$ 4.08 $\rightarrow$ 12.4 1.41 $\rightarrow$ 4.61 $\rightarrow$ 13.5 2.26 $\rightarrow$ 5.63 $\rightarrow$ 18.49 1.24 $\rightarrow$ 4.39 $\rightarrow$ 13.72 1.35 $\rightarrow$ 4.5 $\rightarrow$ 13.84 1.68 $\rightarrow$ 5.52 $\rightarrow$ 15.01 0.86 $\rightarrow$ 3.44 $\rightarrow$ 11.72 1.25 $\rightarrow$ 4.08 $\rightarrow$ 12.25 1.21 $\rightarrow$ 4.12 $\rightarrow$ 12.48 2.27 $\rightarrow$ 5.82 $\rightarrow$ 15.39
+ResNet50 1.68 $\rightarrow$ 3.56 $\rightarrow$ 11.24 2.2 $\rightarrow$ 4.43 $\rightarrow$ 13.11 2.81 $\rightarrow$ 5.62 $\rightarrow$ 17.95 1.89 $\rightarrow$ 4.37 $\rightarrow$ 13.71 1.72 $\rightarrow$ 4.5 $\rightarrow$ 14.21 2.68 $\rightarrow$ 5.18 $\rightarrow$ 14.46 1.02 $\rightarrow$ 2.59 $\rightarrow$ 9.65 1.62 $\rightarrow$ 3.68 $\rightarrow$ 11.27 1.69 $\rightarrow$ 3.68 $\rightarrow$ 11.46 2.92 $\rightarrow$ 5.59 $\rightarrow$ 15.45
+WideResNet50.2 1.34 $\rightarrow$ 3.82 $\rightarrow$ 10.02 1.99 $\rightarrow$ 4.6 $\rightarrow$ 12.09 2.78 $\rightarrow$ 6.06 $\rightarrow$ 17.27 1.48 $\rightarrow$ 4.52 $\rightarrow$ 12.5 1.59 $\rightarrow$ 4.66 $\rightarrow$ 13.02 2.12 $\rightarrow$ 5.13 $\rightarrow$ 12.94 0.93 $\rightarrow$ 2.62 $\rightarrow$ 7.95 1.43 $\rightarrow$ 3.79 $\rightarrow$ 9.96 1.49 $\rightarrow$ 3.67 $\rightarrow$ 10.52 2.58 $\rightarrow$ 5.39 $\rightarrow$ 13.92
+ViT-S+ConvStem 0.89 $\rightarrow$ 2.36 $\rightarrow$ 6.26 1.41 $\rightarrow$ 2.98 $\rightarrow$ 7.76 1.69 $\rightarrow$ 3.34 $\rightarrow$ 10.38 0.97 $\rightarrow$ 2.73 $\rightarrow$ 8.37 1.0 $\rightarrow$ 2.75 $\rightarrow$ 8.51 1.51 $\rightarrow$ 3.09 $\rightarrow$ 8.43 0.32 $\rightarrow$ 1.31 $\rightarrow$ 4.56 0.95 $\rightarrow$ 2.28 $\rightarrow$ 6.37 0.98 $\rightarrow$ 2.31 $\rightarrow$ 6.59 1.53 $\rightarrow$ 3.11 $\rightarrow$ 8.93
+ConvNeXt-B 0.63 $\rightarrow$ 1.56 $\rightarrow$ 4.2 0.9 $\rightarrow$ 1.85 $\rightarrow$ 4.88 0.97 $\rightarrow$ 2.35 $\rightarrow$ 7.5 0.74 $\rightarrow$ 1.85 $\rightarrow$ 5.81 0.74 $\rightarrow$ 1.82 $\rightarrow$ 5.83 1.04 $\rightarrow$ 2.51 $\rightarrow$ 5.6 0.19 $\rightarrow$ 0.74 $\rightarrow$ 2.53 0.69 $\rightarrow$ 1.5 $\rightarrow$ 4.3 0.63 $\rightarrow$ 1.48 $\rightarrow$ 4.22 0.92 $\rightarrow$ 2.48 $\rightarrow$ 6.1
+Swin-B 0.37 $\rightarrow$ 1.35 $\rightarrow$ 3.72 0.63 $\rightarrow$ 1.59 $\rightarrow$ 4.3 0.79 $\rightarrow$ 2.15 $\rightarrow$ 6.93 0.41 $\rightarrow$ 1.62 $\rightarrow$ 5.18 0.45 $\rightarrow$ 1.7 $\rightarrow$ 5.2 0.72 $\rightarrow$ 2.2 $\rightarrow$ 5.55 0.16 $\rightarrow$ 0.72 $\rightarrow$ 2.81 0.37 $\rightarrow$ 1.33 $\rightarrow$ 3.74 0.46 $\rightarrow$ 1.25 $\rightarrow$ 3.95 0.91 $\rightarrow$ 1.65 $\rightarrow$ 5.31
+ConvNeXt-L 0.49 $\rightarrow$ 1.5 $\rightarrow$ 4.35 0.74 $\rightarrow$ 2.07 $\rightarrow$ 5.18 1.07 $\rightarrow$ 2.57 $\rightarrow$ 7.39 0.58 $\rightarrow$ 1.97 $\rightarrow$ 6.45 0.49 $\rightarrow$ 2.07 $\rightarrow$ 6.56 0.74 $\rightarrow$ 2.41 $\rightarrow$ 6.25 0.18 $\rightarrow$ 0.78 $\rightarrow$ 2.31 0.49 $\rightarrow$ 1.61 $\rightarrow$ 4.28 0.49 $\rightarrow$ 1.37 $\rightarrow$ 4.64 1.07 $\rightarrow$ 2.44 $\rightarrow$ 5.93
+ConvV2-L+Swin-L 2.03 $\rightarrow$ 3.46 $\rightarrow$ 6.23 2.65 $\rightarrow$ 4.75 $\rightarrow$ 7.67 1.76 $\rightarrow$ 3.51 $\rightarrow$ 7.52 2.4 $\rightarrow$ 4.38 $\rightarrow$ 9.37 2.48 $\rightarrow$ 4.85 $\rightarrow$ 9.52 2.55 $\rightarrow$ 4.58 $\rightarrow$ 8.83 0.58 $\rightarrow$ 1.51 $\rightarrow$ 3.44 2.0 $\rightarrow$ 3.61 $\rightarrow$ 6.38 1.78 $\rightarrow$ 3.69 $\rightarrow$ 6.9 2.57 $\rightarrow$ 4.8 $\rightarrow$ 9.15
+Swin-L 0.56 $\rightarrow$ 1.25 $\rightarrow$ 3.43 0.85 $\rightarrow$ 1.84 $\rightarrow$ 4.46 1.14 $\rightarrow$ 2.43 $\rightarrow$ 6.84 0.55 $\rightarrow$ 1.51 $\rightarrow$ 5.07 0.64 $\rightarrow$ 1.74 $\rightarrow$ 5.28 0.87 $\rightarrow$ 2.02 $\rightarrow$ 5.12 0.21 $\rightarrow$ 0.59 $\rightarrow$ 2.48 0.56 $\rightarrow$ 1.31 $\rightarrow$ 3.41 0.49 $\rightarrow$ 1.18 $\rightarrow$ 3.61 1.03 $\rightarrow$ 1.92 $\rightarrow$ 4.84
+TABLE XVI: Transfer Attack Performance Across Different Iteration Counts N, 2N, 5N For Epsilon 16/255.
+Model MI-FGSM DI-FGSM TI-FGSM VMI-FGSM VNI-FGSM ADMIX UAP SGM GHOST LGV
+ResNet18 12.4 $\rightarrow$ 13.38 $\rightarrow$ 14.63 13.5 $\rightarrow$ 15.12 $\rightarrow$ 15.77 18.49 $\rightarrow$ 20.23 $\rightarrow$ 20.76 13.72 $\rightarrow$ 15.65 $\rightarrow$ 16.33 13.84 $\rightarrow$ 15.77 $\rightarrow$ 16.22 15.01 $\rightarrow$ 15.24 $\rightarrow$ 16.22 11.72 $\rightarrow$ 11.72 $\rightarrow$ 11.72 12.25 $\rightarrow$ 13.57 $\rightarrow$ 14.37 12.48 $\rightarrow$ 14.14 $\rightarrow$ 15.16 15.39 $\rightarrow$ 17.05 $\rightarrow$ 17.47
+ResNet50 11.24 $\rightarrow$ 12.11 $\rightarrow$ 13.27 13.11 $\rightarrow$ 13.92 $\rightarrow$ 14.89 17.95 $\rightarrow$ 19.29 $\rightarrow$ 20.32 13.71 $\rightarrow$ 14.99 $\rightarrow$ 15.24 14.21 $\rightarrow$ 15.39 $\rightarrow$ 15.52 14.46 $\rightarrow$ 15.08 $\rightarrow$ 15.24 9.65 $\rightarrow$ 9.65 $\rightarrow$ 9.65 11.27 $\rightarrow$ 12.46 $\rightarrow$ 13.18 11.46 $\rightarrow$ 12.86 $\rightarrow$ 14.02 15.45 $\rightarrow$ 17.2 $\rightarrow$ 16.98
+WideResNet50.2 10.02 $\rightarrow$ 10.43 $\rightarrow$ 11.07 12.09 $\rightarrow$ 12.7 $\rightarrow$ 13.08 17.27 $\rightarrow$ 18.88 $\rightarrow$ 19.11 12.5 $\rightarrow$ 13.46 $\rightarrow$ 12.88 13.02 $\rightarrow$ 13.87 $\rightarrow$ 13.55 12.94 $\rightarrow$ 12.38 $\rightarrow$ 12.5 7.95 $\rightarrow$ 7.95 $\rightarrow$ 7.95 9.96 $\rightarrow$ 10.66 $\rightarrow$ 11.07 10.52 $\rightarrow$ 11.1 $\rightarrow$ 11.56 13.92 $\rightarrow$ 14.24 $\rightarrow$ 14.39
+ViT-S+ConvStem 6.26 $\rightarrow$ 6.59 $\rightarrow$ 6.73 7.76 $\rightarrow$ 8.32 $\rightarrow$ 8.15 10.38 $\rightarrow$ 10.38 $\rightarrow$ 10.54 8.37 $\rightarrow$ 9.01 $\rightarrow$ 8.46 8.51 $\rightarrow$ 8.93 $\rightarrow$ 8.6 8.43 $\rightarrow$ 8.43 $\rightarrow$ 8.09 4.56 $\rightarrow$ 4.56 $\rightarrow$ 4.56 6.37 $\rightarrow$ 6.68 $\rightarrow$ 6.87 6.59 $\rightarrow$ 6.82 $\rightarrow$ 7.04 8.93 $\rightarrow$ 9.18 $\rightarrow$ 8.54
+ConvNeXt-B 4.2 $\rightarrow$ 4.12 $\rightarrow$ 4.3 4.88 $\rightarrow$ 5.17 $\rightarrow$ 5.12 7.5 $\rightarrow$ 8.47 $\rightarrow$ 8.13 5.81 $\rightarrow$ 6.04 $\rightarrow$ 5.54 5.83 $\rightarrow$ 5.94 $\rightarrow$ 5.49 5.6 $\rightarrow$ 5.3 $\rightarrow$ 5.28 2.53 $\rightarrow$ 2.53 $\rightarrow$ 2.53 4.3 $\rightarrow$ 4.28 $\rightarrow$ 4.22 4.22 $\rightarrow$ 4.35 $\rightarrow$ 4.57 6.1 $\rightarrow$ 6.04 $\rightarrow$ 6.12
+Swin-B 3.72 $\rightarrow$ 3.9 $\rightarrow$ 4.06 4.3 $\rightarrow$ 4.43 $\rightarrow$ 5.18 6.93 $\rightarrow$ 7.3 $\rightarrow$ 7.4 5.18 $\rightarrow$ 5.36 $\rightarrow$ 5.2 5.2 $\rightarrow$ 5.57 $\rightarrow$ 5.33 5.55 $\rightarrow$ 5.12 $\rightarrow$ 4.94 2.81 $\rightarrow$ 2.81 $\rightarrow$ 2.81 3.74 $\rightarrow$ 3.95 $\rightarrow$ 4.03 3.95 $\rightarrow$ 4.01 $\rightarrow$ 4.22 5.31 $\rightarrow$ 5.6 $\rightarrow$ 5.65
+ConvNeXt-L 4.35 $\rightarrow$ 4.02 $\rightarrow$ 4.09 5.18 $\rightarrow$ 5.13 $\rightarrow$ 5.52 7.39 $\rightarrow$ 7.62 $\rightarrow$ 7.98 6.45 $\rightarrow$ 6.14 $\rightarrow$ 5.52 6.56 $\rightarrow$ 6.4 $\rightarrow$ 5.7 6.25 $\rightarrow$ 5.62 $\rightarrow$ 5.6 2.31 $\rightarrow$ 2.31 $\rightarrow$ 2.31 4.28 $\rightarrow$ 4.02 $\rightarrow$ 4.15 4.64 $\rightarrow$ 4.2 $\rightarrow$ 4.41 5.93 $\rightarrow$ 6.12 $\rightarrow$ 5.62
+ConvV2-L+Swin-L 6.23 $\rightarrow$ 6.41 $\rightarrow$ 6.51 7.67 $\rightarrow$ 8.31 $\rightarrow$ 8.43 7.52 $\rightarrow$ 7.99 $\rightarrow$ 8.06 9.37 $\rightarrow$ 9.2 $\rightarrow$ 8.61 9.52 $\rightarrow$ 9.55 $\rightarrow$ 8.98 8.83 $\rightarrow$ 8.34 $\rightarrow$ 8.19 3.44 $\rightarrow$ 3.44 $\rightarrow$ 3.44 6.38 $\rightarrow$ 6.36 $\rightarrow$ 6.6 6.9 $\rightarrow$ 6.73 $\rightarrow$ 6.73 9.15 $\rightarrow$ 9.03 $\rightarrow$ 8.9
+Swin-L 3.43 $\rightarrow$ 3.53 $\rightarrow$ 3.61 4.46 $\rightarrow$ 4.38 $\rightarrow$ 4.51 6.84 $\rightarrow$ 7.32 $\rightarrow$ 6.97 5.07 $\rightarrow$ 4.94 $\rightarrow$ 4.64 5.28 $\rightarrow$ 5.25 $\rightarrow$ 4.92 5.12 $\rightarrow$ 4.53 $\rightarrow$ 4.43 2.48 $\rightarrow$ 2.48 $\rightarrow$ 2.48 3.41 $\rightarrow$ 3.51 $\rightarrow$ 3.53 3.61 $\rightarrow$ 3.51 $\rightarrow$ 3.53 4.84 $\rightarrow$ 4.89 $\rightarrow$ 4.92
+TABLE XVII: Query Attack Performance Across Increased Epsilon 4/255, 8/255 and 16/255.
+Model BASES TREMBA
+ResNet18 2.36 $\rightarrow$ 5.33 $\rightarrow$ 13.43 4.06 $\rightarrow$ 11.3 $\rightarrow$ 37.39
+ResNet50 2.83 $\rightarrow$ 5.09 $\rightarrow$ 13.76 3.56 $\rightarrow$ 9.52 $\rightarrow$ 32.67
+WideResNet50.2 2.35 $\rightarrow$ 5.27 $\rightarrow$ 12.07 3.22 $\rightarrow$ 9.04 $\rightarrow$ 31.77
+ViT-S+ConvStem 1.55 $\rightarrow$ 3.39 $\rightarrow$ 7.93 1.54 $\rightarrow$ 5.42 $\rightarrow$ 24.11
+ConvNeXt-B 1.38 $\rightarrow$ 2.53 $\rightarrow$ 5.8 1.34 $\rightarrow$ 4.46 $\rightarrow$ 18.54
+Swin-B 0.96 $\rightarrow$ 2.26 $\rightarrow$ 5.2 1.14 $\rightarrow$ 4.11 $\rightarrow$ 17.17
+ConvNeXt-L 1.16 $\rightarrow$ 2.7 $\rightarrow$ 6.04 1.47 $\rightarrow$ 4.43 $\rightarrow$ 17.21
+ConvV2-L+Swin-L 4.77 $\rightarrow$ 6.73 $\rightarrow$ 10.12 2.94 $\rightarrow$ 7.59 $\rightarrow$ 21.84
+Swin-L 1.26 $\rightarrow$ 2.18 $\rightarrow$ 4.79 1.25 $\rightarrow$ 4.02 $\rightarrow$ 16.41
+TABLE XVIII: Query Attack Performance Across Different Query Counts N, 2N For Epsilon 16/255.
+Model BASES TREMBA
+ResNet18 13.43 $\rightarrow$ 13.73 37.39 $\rightarrow$ 43.06
+ResNet50 13.76 $\rightarrow$ 14.05 32.67 $\rightarrow$ 37.78
+WideResNet50.2 12.07 $\rightarrow$ 12.23 31.77 $\rightarrow$ 36.24
+ViT-S+ConvStem 7.93 $\rightarrow$ 8.07 24.11 $\rightarrow$ 29.45
+ConvNeXt-B 5.8 $\rightarrow$ 5.93 18.54 $\rightarrow$ 22.51
+Swin-B 5.2 $\rightarrow$ 5.2 17.17 $\rightarrow$ 21.44
+ConvNeXt-L 6.04 $\rightarrow$ 6.17 17.21 $\rightarrow$ 21.02
+ConvV2-L+Swin-L 10.12 $\rightarrow$ 10.29 21.84 $\rightarrow$ 26.34
+Swin-L 4.79 $\rightarrow$ 4.99 16.41 $\rightarrow$ 20.46
+Figure 8: Success rate of blackbox attacks against SoTA defenses for epsillon 8/255.
+Figure 9: Success rate of blackbox attacks against SoTA defenses for epsilon 16/255.
+Figure 10: Success rate of blackbox attacks against SoTA defenses for epsilon 16/255 and iteration=2N.
+Figure 11: Success rate of blackbox attacks against SoTA defenses for epsilon 16/255 and iteration=5N.
+B-A AutoAttack Robustness of LGV and GHOST Models
 To evaluate the impact of the LGV weight collection phase, and the GHOST perturbing phase on model robustness, we collected 10 models after finishing the first phase for both LGV and GHOST, prior to proceeding to the attacking phase.
 Specifically, the LGV variations were obtained by training the surrogate model for additional epochs using a high learning rate, during which the weights of the models were collected periodically.
-For GHOST networks, the variations were obtained by applying random perturbations to the skip connections of the surrogate model.
-Using the robust surrogate model WideResNet-101-2 introduced in (Peng et al., 2023) and described in Section 5, we evaluated the robustness of the resulting LGV models and compared it with the robustness of GHOST models.
-The robust accuracies for the collected models are reported in Table 9.
+For GHOST networks, the variations were obtained by introducing a random perturbation to the scaling factor in the skip connections.
+Using the robust surrogate model WideResNet-101-2 introduced in [54] and described in Section V, we evaluated the robustness of the resulting LGV models and compared it with the robustness of GHOST models.
+The robust accuracies for the collected models are reported in Table XIX.
 Given that both LGV and GHOST models are derived from the same robust surrogate model, we observed a difference in their robustness.
@@
 This ensures that any preliminary training phase does not inadvertently weaken the surrogate robustness.
-Table 9: Robust accuracies of LGV and GHOST models before the attacking phase.
+TABLE XIX: Robust accuracies of LGV and GHOST models before the attacking phase.
 Num 1 Num 2 Num 3 Num 4 Num 5 Num 6 Num 7 Num 8 Num 9 Num 10 Average
@@
 LGV 0 0 0 0 0 0 0 0 0 0 0
-Appendix E Detailed Tabular Results
+B-B Detailed Tabular Results
 We provide the tabular representation corresponding to the results presented in the main paper.
-Tables 10, 11, 12, 13 represent the success rates for section 3, 4, and 5 with vanilla surrogates and robust surrogates, respectively.
+Tables XX, XXI, XXII, XXIII represent the success rates for section III, IV, and V with vanilla surrogates and robust surrogates, respectively.
 These tables display the mean and standard deviation of success rates computed across three random seeds.
-Table 10: Success rates of black-box attacks against one vanilla, and one robust ResNet50 model.
+B-C Reproduction Package and Availability
+The source code to reproduce the experiments of this paper is available at https://github.com/serval-uni-lu/RobustBlack.git.
+TABLE $X$: Success rates of black-box attacks against one vanilla, and one robust ResNet50 model.
 MODELS MI-FGSM DI-FGSM TI-FGSM VMI-FGSM VNI-FGSM ADMIX UAP GHOST LGV BASES TREMBA
-Vanilla ResNet50 57.77 $\pm$0.00 71.31 $\pm$0.32 43.59 $\pm$ 0.17 66.79 $\pm$ 0.07 68.00 $\pm$ 0.18 99.95 $\pm$ 0.00 8.89 $\pm$ 0.20 64.53 $\pm$ 0.13 19.81 $\pm$ 1.46 81.08 $\pm$0.49 89.56 $\pm$0.09
-Robust ResNet50 1.53 $\pm$ 0.00 1.94 $\pm$ 0.03 2.3 $\pm$ 0.01 1.66 $\pm$ 0.03 1.85 $\pm$ 0.04 2.54 $\pm$ 0.09 0.98 $\pm$ 0.10 1.68 $\pm$0.03 1.61 $\pm$0.10 2.83 $\pm$0.04 3.56 $\pm$0.06
-Table 11: Success rates of white-box AutoAttack and black-box attacks against robust targets from Robustbench.
+Vanilla ResNet50 57.77 $\pm$0.00 71.31 $\pm$0.32 43.59 $\pm$ 0.17 66.79 $\pm$ 0.07 68.00 $\pm$ 0.18 99.95 $\pm$ 0.00 8.89 $\pm$ 0.20 64.53 $\pm$ 0.13 90.11 $\pm$ 0.12 81.08 $\pm$0.49 89.56 $\pm$0.09
+Robust ResNet50 1.53 $\pm$ 0.00 1.94 $\pm$ 0.03 2.3 $\pm$ 0.01 1.66 $\pm$ 0.03 1.85 $\pm$ 0.04 2.54 $\pm$ 0.09 0.98 $\pm$ 0.10 1.68 $\pm$0.03 1.33 $\pm$0.01 2.83 $\pm$0.04 3.56 $\pm$0.06
+TABLE XXI: Success rates of white-box AutoAttack and black-box attacks against robust targets from Robustbench.
 RANK MODELS AUTOATTACK MI-FGSM DI-FGSM TI-FGSM VMI-FGSM VNI-FGSM ADMIX UAP SGM GHOST LGV BASES TREMBA
-21 Salman ResNet18 50.53 1.28 $\pm$0.00 1.41 $\pm$0.02 2.26 $\pm$ 0.04 1.24 $\pm$ 0.05 1.35 $\pm$ 0.02 1.68 $\pm$ 0.06 0.86 $\pm$ 0.10 1.25 $\pm$ 0.00 1.21 $\pm$0.18 1.63 $\pm$0.13 2.36 $\pm$0.04 4.06 $\pm$0.06
-18 Salman ResNet50 43.84 1.68 $\pm$ 0.00 2.2 $\pm$ 0.10 2.81 $\pm$ 0.07 1.89 $\pm$ 0.04 1.72 $\pm$ 0.03 2.68 $\pm$ 0.13 1.02 $\pm$ 0.01 1.62 $\pm$ 0.00 1.69 $\pm$0.05 1.72 $\pm$0.26 2.83 $\pm$0.04 3.56 $\pm$0.06
-17 Salman WideResNet50.2 42.87 1.34 $\pm$ 0.00 1.99 $\pm$ 0.04 2.78 $\pm$ 0.10 1.48 $\pm$ 0.05 1.59 $\pm$ 0.05 2.12 $\pm$ 0.08 0.93 $\pm$ 0.07 1.43 $\pm$ 0.00 1.49 $\pm$0.06 1.52 $\pm$0.08 2.35 $\pm$0.06 3.22 $\pm$0.06
-12 Sing ViT-S, ConvStem 34.76 0.89 $\pm$ 0.00 1.41 $\pm$ 0.05 1.69 $\pm$ 0.05 0.97 $\pm$ 0.04 1.00 $\pm$ 0.00 1.51 $\pm$ 0.05 0.32 $\pm$ 0.06 0.95 $\pm$ 0.00 0.98 $\pm$0.03 0.72 $\pm$0.12 1.55 $\pm$0.00 1.54 $\pm$0.03
-7 Liu ConvNeXt-B 28.01 0.63 $\pm$ 0.00 0.9 $\pm$ 0.02 0.97 $\pm$ 0.02 0.74 $\pm$ 0.02 0.74 $\pm$ 0.04 1.04 $\pm$ 0.04 0.19 $\pm$ 0.04 0.69 $\pm$ 0.00 0.63 $\pm$0.03 0.57 $\pm$0.04 1.38 $\pm$0.06 1.34 $\pm$0.04
-5 Liu Swin-B 27.14 0.37 $\pm$ 0.00 0.63 $\pm$ 0.03 0.79 $\pm$ 0.07 0.41 $\pm$ 0.01 0.45 $\pm$ 0.02 0.72 $\pm$ 0.02 0.16 $\pm$ 0.02 0.37 $\pm$ 0.00 0.46 $\pm$0.07 0.36 $\pm$0.05 0.96 $\pm$0.02 1.14 $\pm$0.03
-3 Liu ConvNeXt-L 25.95 0.49 $\pm$ 0.00 0.74 $\pm$ 0.09 1.07 $\pm$ 0.03 0.58 $\pm$ 0.03 0.49 $\pm$ 0.02 0.74 $\pm$ 0.01 0.18 $\pm$ 0.06 0.49 $\pm$ 0.00 0.49 $\pm$0.09 0.45 $\pm$0.05 1.16 $\pm$0.00 1.47 $\pm$0.01
-2 Bai ConvNeXtV2-L, Swin-L 25.48 2.03 $\pm$ 0.00 2.65 $\pm$ 0.03 1.76 $\pm$ 0.05 2.40 $\pm$ 0.05 2.48 $\pm$ 0.03 2.55 $\pm$ 0.05 0.58 $\pm$ 0.05 2.00 $\pm$ 0.00 1.78 $\pm$0.16 1.01 $\pm$0.14 4.77 $\pm$0.13 2.94 $\pm$0.12
-1 Liu Swin-L 25.98 0.56 $\pm$ 0.00 0.85 $\pm$ 0.04 1.14 $\pm$ 0.08 0.55 $\pm$ 0.01 0.64 $\pm$ 0.04 0.87 $\pm$ 0.00 0.21 $\pm$ 0.04 0.56 $\pm$ 0.00 0.49 $\pm$0.04 0.46 $\pm$0.04 1.26 $\pm$0.02 1.25 $\pm$0.01
-Table 12: Success rates of black-box attacks using vanilla surrogates against a vanilla target and robust targets.
-Models GHOST LGV BASES TREMBA VMI-FGSM TI-FGSM ADMIX VNI-FGSM DI-FGSM UAP MI-FGSM
-Standard ResNet50 49.36 $\pm$ 0.14 15.98 $\pm$ 1.13 81.08 $\pm$ 0.49 89.56 $\pm$ 0.09 61.22 $\pm$ 0.18 41.5 $\pm$ 0.35 64.75 $\pm$ 0.17 61.10 $\pm$ 0.13 65.95 $\pm$ 0.75 7.11 $\pm$ 0.25 51.62 $\pm$ 0.0
-Salman ResNet18 1.13 $\pm$ 0.08 1.37 $\pm$ 0.12 2.36 $\pm$ 0.04 4.06 $\pm$ 0.06 1.29 $\pm$ 0.05 2.04 $\pm$ 0.06 1.88 $\pm$ 0.08 1.35 $\pm$ 0.06 1.51 $\pm$ 0.03 0.83 $\pm$ 0.03 0.98 $\pm$ 0.0
-Liu ConvNeXt-L 0.48 $\pm$ 0.09 0.45 $\pm$ 0.01 1.16 $\pm$ 0.0 1.47 $\pm$ 0.01 0.7 $\pm$ 0.02 1.19 $\pm$ 0.1 0.99 $\pm$ 0.03 0.8 $\pm$ 0.04 0.92 $\pm$ 0.05 0.18 $\pm$ 0.04 0.6 $\pm$ 0.0
-Liu ConvNeXtV2-L+Swin-L 1.7 $\pm$ 0.19 1.0 $\pm$ 0.03 4.77 $\pm$ 0.13 2.94 $\pm$ 0.12 2.51 $\pm$ 0.02 2.08 $\pm$ 0.04 2.52 $\pm$ 0.06 2.45 $\pm$ 0.05 2.79 $\pm$ 0.07 0.63 $\pm$ 0.03 2.0 $\pm$ 0.0
-Liu Swin-L 0.53 $\pm$ 0.03 0.45 $\pm$ 0.01 1.26 $\pm$ 0.02 1.25 $\pm$ 0.01 0.74 $\pm$ 0.02 1.22 $\pm$ 0.05 0.93 $\pm$ 0.05 0.74 $\pm$ 0.04 0.88 $\pm$ 0.03 0.18 $\pm$ 0.0 0.51 $\pm$ 0.0
-Table 13: Success rates of black-box attacks using robust surrogates against a vanilla target and robust targets.
-Models GHOST LGV BASES TREMBA VMI-FGSM TI-FGSM ADMIX VNI-FGSM DI-FGSM UAP MI-FGSM
-Standard ResNet50 17.95 $\pm$ 0.36 7.79 $\pm$ 1.58 17.94 $\pm$ 0.00 67.54 $\pm$ 0.13 16.47 $\pm$ 0.02 8.48 $\pm$ 0.18 15.22 $\pm$ 0.11 16.51 $\pm$ 0.04 15.28 $\pm$ 0.21 0.73 $\pm$ 0.0 15.96 $\pm$ 0.0
-Salman ResNet18 14.15 $\pm$ 0.22 2.04 $\pm$ 0.19 11.95 $\pm$ 0.00 10.58 $\pm$ 0.14 12.54 $\pm$ 0.02 12.39 $\pm$ 0.04 14.4 $\pm$ 0.17 12.34 $\pm$ 0.05 13.83 $\pm$ 0.14 0.79 $\pm$ 0.0 12.44 $\pm$ 0.0
-Liu ConvNeXt-L 9.23 $\pm$ 0.13 0.42 $\pm$ 0.12 12.02 $\pm$ 0.00 4.94 $\pm$ 0.06 9.71 $\pm$ 0.01 7.28 $\pm$ 0.04 9.1 $\pm$ 0.06 9.46 $\pm$ 0.00 9.74 $\pm$ 0.0 0.36 $\pm$ 0.0 9.67 $\pm$ 0.0
-Liu ConvNeXtV2-L+Swin-L 6.69 $\pm$ 0.08 0.88 $\pm$ 0.15 11.77 $\pm$ 0.01 5.33 $\pm$ 0.11 7.54 $\pm$ 0.03 5.09 $\pm$ 0.04 7.0 $\pm$ 0.02 7.57 $\pm$ 0.00 7.07 $\pm$ 0.09 0.37 $\pm$ 0.0 7.59 $\pm$ 0.0
-Liu Swin-L 8.16 $\pm$ 0.09 0.55 $\pm$ 0.13 12.55 $\pm$ 0.00 4.4 $\pm$ 0.08 8.95 $\pm$ 0.01 6.78 $\pm$ 0.07 8.21 $\pm$ 0.01 8.89 $\pm$ 0.02 8.72 $\pm$ 0.04 0.31 $\pm$ 0.0 8.94 $\pm$ 0.0
+21 Salman ResNet18 50.53 1.28 $\pm$0.00 1.41 $\pm$0.02 2.26 $\pm$ 0.04 1.24 $\pm$ 0.05 1.35 $\pm$ 0.02 1.68 $\pm$ 0.06 0.86 $\pm$ 0.10 1.25 $\pm$ 0.00 1.21 $\pm$0.18 2.27 $\pm$0.11 2.36 $\pm$0.04 4.06 $\pm$0.06
+18 Salman ResNet50 43.84 1.68 $\pm$ 0.00 2.2 $\pm$ 0.10 2.81 $\pm$ 0.07 1.89 $\pm$ 0.04 1.72 $\pm$ 0.03 2.68 $\pm$ 0.13 1.02 $\pm$ 0.01 1.62 $\pm$ 0.00 1.69 $\pm$0.05 2.92 $\pm$0.01 2.83 $\pm$0.04 3.56 $\pm$0.06
+17 Salman WideResNet50.2 42.87 1.34 $\pm$ 0.00 1.99 $\pm$ 0.04 2.78 $\pm$ 0.10 1.48 $\pm$ 0.05 1.59 $\pm$ 0.05 2.12 $\pm$ 0.08 0.93 $\pm$ 0.07 1.43 $\pm$ 0.00 1.49 $\pm$0.06 2.58 $\pm$0.09 2.35 $\pm$0.06 3.22 $\pm$0.06
+12 Sing ViT-S, ConvStem 34.76 0.89 $\pm$ 0.00 1.41 $\pm$ 0.05 1.69 $\pm$ 0.05 0.97 $\pm$ 0.04 1.00 $\pm$ 0.00 1.51 $\pm$ 0.05 0.32 $\pm$ 0.06 0.95 $\pm$ 0.00 0.98 $\pm$0.03 1.53 $\pm$0.08 1.55 $\pm$0.00 1.54 $\pm$0.03
+7 Liu ConvNeXt-B 28.01 0.63 $\pm$ 0.00 0.9 $\pm$ 0.02 0.97 $\pm$ 0.02 0.74 $\pm$ 0.02 0.74 $\pm$ 0.04 1.04 $\pm$ 0.04 0.19 $\pm$ 0.04 0.69 $\pm$ 0.00 0.63 $\pm$0.03 0.92 $\pm$0.07 1.38 $\pm$0.06 1.34 $\pm$0.04
+5 Liu Swin-B 27.14 0.37 $\pm$ 0.00 0.63 $\pm$ 0.03 0.79 $\pm$ 0.07 0.41 $\pm$ 0.01 0.45 $\pm$ 0.02 0.72 $\pm$ 0.02 0.16 $\pm$ 0.02 0.37 $\pm$ 0.00 0.46 $\pm$0.07 0.91 $\pm$0.03 0.96 $\pm$0.02 1.14 $\pm$0.03
+3 Liu ConvNeXt-L 25.95 0.49 $\pm$ 0.00 0.74 $\pm$ 0.09 1.07 $\pm$ 0.03 0.58 $\pm$ 0.03 0.49 $\pm$ 0.02 0.74 $\pm$ 0.01 0.18 $\pm$ 0.06 0.49 $\pm$ 0.00 0.49 $\pm$0.09 1.07 $\pm$0.03 1.16 $\pm$0.00 1.47 $\pm$0.01
+2 Bai ConvNeXtV2-L, Swin-L 25.48 2.03 $\pm$ 0.00 2.65 $\pm$ 0.03 1.76 $\pm$ 0.05 2.40 $\pm$ 0.05 2.48 $\pm$ 0.03 2.55 $\pm$ 0.05 0.58 $\pm$ 0.05 2.00 $\pm$ 0.00 1.78 $\pm$0.16 2.57 $\pm$0.10 4.77 $\pm$0.13 2.94 $\pm$0.12
+1 Liu Swin-L 25.98 0.56 $\pm$ 0.00 0.85 $\pm$ 0.04 1.14 $\pm$ 0.08 0.55 $\pm$ 0.01 0.64 $\pm$ 0.04 0.87 $\pm$ 0.00 0.21 $\pm$ 0.04 0.56 $\pm$ 0.00 0.49 $\pm$0.04 1.03 $\pm$0.03 1.26 $\pm$0.02 1.25 $\pm$0.01
+TABLE XXII: Success rates of black-box attacks using vanilla surrogates against a vanilla target and robust targets.
+Models MI-FGSM DI-FGSM TI-FGSM VMI-FGSM VNI-FGSM ADMIX UAP GHOST LGV BASES TREMBA
+Standard ResNet50 51.62 $\pm$ 0.0 65.95 $\pm$ 0.75 41.5 $\pm$ 0.35 61.22 $\pm$ 0.18 61.10 $\pm$ 0.13 64.75 $\pm$ 0.17 7.11 $\pm$ 0.25 49.36 $\pm$ 0.14 87.39 $\pm$ 0.26 81.08 $\pm$ 0.49 89.56 $\pm$ 0.09
+Salman ResNet18 0.98 $\pm$ 0.0 1.51 $\pm$ 0.03 2.04 $\pm$ 0.06 1.29 $\pm$ 0.05 1.35 $\pm$ 0.06 1.88 $\pm$ 0.08 0.83 $\pm$ 0.03 1.13 $\pm$ 0.08 2.17 $\pm$ 0.05 2.36 $\pm$ 0.04 4.06 $\pm$ 0.06
+Liu ConvNeXt-L 0.6 $\pm$ 0.0 0.92 $\pm$ 0.05 1.19 $\pm$ 0.1 0.7 $\pm$ 0.02 0.8 $\pm$ 0.04 0.99 $\pm$ 0.03 0.18 $\pm$ 0.04 0.48 $\pm$ 0.09 1.05 $\pm$ 0.02 1.16 $\pm$ 0.0 1.47 $\pm$ 0.01
+Liu ConvNeXtV2-L+Swin-L 2.0 $\pm$ 0.0 2.79 $\pm$ 0.07 2.08 $\pm$ 0.04 2.51 $\pm$ 0.02 2.45 $\pm$ 0.05 2.52 $\pm$ 0.06 0.63 $\pm$ 0.03 1.7 $\pm$ 0.19 2.57 $\pm$ 0.07 4.77 $\pm$ 0.13 2.94 $\pm$ 0.12
+Liu Swin-L 0.51 $\pm$ 0.0 0.88 $\pm$ 0.03 1.22 $\pm$ 0.05 0.74 $\pm$ 0.02 0.74 $\pm$ 0.04 0.93 $\pm$ 0.05 0.18 $\pm$ 0.0 0.53 $\pm$ 0.03 1.1 $\pm$ 0.11 1.26 $\pm$ 0.02 1.25 $\pm$ 0.01
+TABLE XXIII: Success rates of black-box attacks using robust surrogates against a vanilla target and robust targets.
+Models MI-FGSM DI-FGSM TI-FGSM VMI-FGSM VNI-FGSM ADMIX UAP GHOST LGV BASES TREMBA
+Standard ResNet50 15.96 $\pm$ 0.0 15.28 $\pm$ 0.21 8.48 $\pm$ 0.18 16.47 $\pm$ 0.02 16.51 $\pm$ 0.04 15.22 $\pm$ 0.11 0.73 $\pm$ 0.0 17.95 $\pm$ 0.36 81.34 $\pm$ 0.43 17.94 $\pm$ 0.00 67.54 $\pm$ 0.13
+Salman ResNet18 12.44 $\pm$ 0.0 13.83 $\pm$ 0.14 12.39 $\pm$ 0.04 12.54 $\pm$ 0.02 12.34 $\pm$ 0.05 14.4 $\pm$ 0.17 0.79 $\pm$ 0.0 14.15 $\pm$ 0.22 2.78 $\pm$ 0.21 11.95 $\pm$ 0.00 10.58 $\pm$ 0.14
+Liu ConvNeXt-L 9.67 $\pm$ 0.0 9.74 $\pm$ 0.0 7.28 $\pm$ 0.04 9.71 $\pm$ 0.01 9.46 $\pm$ 0.00 9.1 $\pm$ 0.06 0.36 $\pm$ 0.0 9.23 $\pm$ 0.13 1.6 $\pm$ 0.11 12.02 $\pm$ 0.00 4.94 $\pm$ 0.06
+Liu ConvNeXtV2-L+Swin-L 7.59 $\pm$ 0.0 7.07 $\pm$ 0.09 5.09 $\pm$ 0.04 7.54 $\pm$ 0.03 7.57 $\pm$ 0.00 7.0 $\pm$ 0.02 0.37 $\pm$ 0.0 6.69 $\pm$ 0.08 4.9 $\pm$ 0.06 11.77 $\pm$ 0.01 5.33 $\pm$ 0.11
+Liu Swin-L 8.94 $\pm$ 0.0 8.72 $\pm$ 0.04 6.78 $\pm$ 0.07 8.95 $\pm$ 0.01 8.89 $\pm$ 0.02 8.21 $\pm$ 0.01 0.31 $\pm$ 0.0 8.16 $\pm$ 0.09 1.57 $\pm$ 0.04 12.55 $\pm$ 0.00 4.4 $\pm$ 0.08